Drawing carbocyclic compounds by carom.sty (v2.00a) of XΥMTEX

Shinsaku Fujita Department of Chemistry and Materials Technology, Kyoto Institute of Technology, Matsugasaki, Sakyoku, Kyoto, 606-8585 Japan

2000/06/14

1 Introduction

1.1 Options for docstrip

option function
carom carom.sty
driver driver for this dtx file

1.2 Version Information

1 ⟨∗carom⟩

2 \typeout{XyMTeX for Drawing Chemical Structural Formulas. Version 2.00}
3 \typeout{       -- Released December 25, 1998 by Shinsaku Fujita}
4 % %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
5 \def\j@urnalname{carom}
6 \def\versi@ndate{June 14, 2000}
7 \def\versi@nno{ver2.00a}
8 \def\copyrighth@lder{SF} % Shinsaku Fujita
9 % %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
10 \typeout{XyMTeX Macro File `\j@urnalname' (\versi@nno) <\versi@ndate>%
11 \space[\copyrighth@lder]}

2 List of commands for carom.sty

*******************************
* carom.sty: list of commands *
*******************************
  <VERTICAL TYPE>
    \cyclohexanev               \@cyclohexanev

\bzdrv                      \@bzdrv
\decalinev                  \@decalinev
\decalinevb                 \@decalinevb
\decalinevt                 \@decalinevt
\naphdrv                    \@naphdrv
\naphdrvb                   \@naphdrvb
\naphdrvt                   \@naphdrvt
\tetralinev                 \@tetralinev
\tetralinevb                \@tetralinevb
\tetralinevt                \@tetralinevt
\hanthracenev               \@hanthracenev
\anthracenev                \@anthracenev
\hphenanthrenev             \@hphenanthrenev
\phenanthrenev              \@phenanthrenev
\steroid                    \@steroid
\steroidchain               \@steroidchain

<HORIZONTAL TYPE>
\cyclohexaneh               \@cyclohexaneh
\bzdrh                      \@bzdrh
\decalineh                  \@decalineh
\naphdrh                    \@naphdrh
\tetralineh                 \@tetralineh

3 Input of basic macros

To assure the compatibility to L^{A}TEX2.09 (the native mode), the commands added by L^{A}TEX 2ε have not been used in the resulting sty files (carom.sty for the present case). Hence, the combination of \input and \@ifundefined is used to crossload sty files (chemstr.sty for the present case) in place of the \RequirePackage command of L^{A}TEX 2ε.

12 % *************************
13 % * input of basic macros *
14 % *************************
15 \@ifundefined{setsixringv}{\input chemstr.sty\relax}{}
16 \@ifundefined{decaheterov}{\input hetarom.sty\relax}{}
17 \@ifundefined{decaheteroh}{\input hetaromh.sty\relax}{}
18 \unitlength=0.1pt

4 Cyclohexane derivatives

4.1 Vertical cyclohexanes

The macro \cyclohexanev is used for drawing cyclohexane derivatives of vertical type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \setsixringv.

 ***************************
 * cyclohexane derivatives *
 *  (vertical type)        *
 ***************************
 The following numbering is adopted in this macro.
          1
          *
     6  *   *  2
       |     |
       |     |
     5  *   *  3
          *
          4 <===== the original point

The macro \cyclohexanev has an argument SUBSLIST as well as an optional argument BONDLIST.

   \cyclohexanev[BONDLIST]{SUBSLIST}

The BONDLIST argument contains one or more characters selected form a to f as well as A, each of which indicates the presence of an inner (endcyclic) double bond on the corresponding position.

     BONDLIST =
           none       :  cyclohexane
           a          :  1,2-double bond
           b          :  2,3-double bond
           c          :  4,3-double bond
           d          :  4,5-double bond
           e          :  5,6-double bond
           f          :  6,1-double bond
           A          :  aromatic circle

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 6
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specifed) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)
                for the other ring fusion
nGB        :  beta single bond at n-atom (boldface)
                for the other ring fusion

e.g.
\cyclohexanev{1==Cl;2==F}
\cyclohexanev[c]{1==Cl;4==F;2==CH$_{3}$}
\cyclohexanev[eb]{1D==O;4SA==MeO;4SB==OMe;5W==Cl;6==Cl}

\cyclohexanev
19 \def\cyclohexanev{\@ifnextchar[{\@cyclohexanev[@}{\@cyclohexanev[@]}}
20 \def\@cyclohexanev[#1]#2{\sixheterov[#1]{}{#2}}

The following definition has been deleted.

(Old defition)
% \begin{macro}{\cyclohexanev}
%    \begin{macrocode}
\def\cyclohexanev{\@ifnextchar[{\@cyclohexanev[@}{\@cyclohexanev[@]}}
\def\@cyclohexanev[#1]#2{%
\iniatom\iniflag%initialize
\@reset@ylsw%

\ylposition{#2}{0}{0}{7}{0}%
\if@ylsw
 \yl@shiftii=\@ylii
 \yl@shifti=\@yli
 \advance\yl@shiftii\yl@xdiff
 \advance\yl@shifti\yl@ydiff
 \begin{picture}(0,0)(-\yl@shiftii,-\yl@shifti)
 \reset@yl@xydiff%1999/1/6 by S. Fujita
\else
 \begin{picture}(800,880)(-\shiftii,-\shifti)
  \iforigpt \put(-\shiftii,-\shifti){\circle*{50}}%
           \put(-\noshift,-\noshift){\circle{50}}%
   \typeout{command `cyclohexanev' origin: %
    (\the\noshift,\the\noshift) ---> (\the\shiftii,\the\shifti)}\fi%
\fi
  \put(0,406){\line(-5,-3){171}}%       %bond 1-6     8-7
  \put(0,406){\line(5,-3){171}}%        %bond 1-2     8-8a
  \put(0,0){\line(-5,3){171}}%          %bond 4-5     5-6
  \put(0,0){\line(5,3){171}}%           %bond 4-3     5-4a
  \put(171,103){\line(0,1){200}}%       %bond 3-2     4a-8a
  \put(-171,103){\line(0,1){200}}%      %bond 5-6     6-7
\@tfor\member:=#1\do{%
\if\member r%no endocyclic double bonds
\else\if\member a%
  \put(6,364){\line(5,-3){126}}%        %double bond 1-2    8-8a
\else\if\member b%
  \put(138,129){\line(0,1){148}}%       %double bond 3-2    4a-8a
\else\if\member c%
  \put(6,42){\line(5,3){126}}%          %double bond 4-3    5-4a
\else\if\member d%
  \put(-6,42){\line(-5,3){126}}%        %double bond 4-5    5-6
\else\if\member e%
  \put(-138,129){\line(0,1){148}}%      %double bond 5-6    6-7
\else\if\member f%
  \put(-6,364){\line(-5,-3){126}}%      %double bond 1-6    8-7
\else\if\member A%left aromatic circle
  \put(0,203){\circle{240}}            %circle
\else \expandafter\twoCH@@R\member//%
      \set@fusionadd@sixv
\fi\fi\fi\fi\fi\fi\fi\fi}%
% %%%%%%%%%%%%%%%%%%%%%%%%
% % setting substituents %
% %%%%%%%%%%%%%%%%%%%%%%%%
\setsixringv{#2}{0}{0}{7}{0}%
\end{picture}}%end of \cyclohexanev macro
%    \end{macrocode}
% \end{macro}

4.2 Horizontal cyclohexanes

The macro cyclohexaneh is used for drawing cyclohexane derivatives of vertical type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \setsixringh.

 ***************************
 * cyclohexane derivatives *
 *  (horizontal type)      *
 ***************************
 The following numbering is adopted in this macro.
                                2      3
                                 -----
                               *       *
   the original point ===> 1 *           * 4
          (0,0)                *       *
                                 -----
                                6      5

The macro \cyclohexaneh has an argument SUBSLIST as well as an optional argument BONDLIST.

   \cyclohexaneh[BONDLIST]{SUBSLIST}

The BONDLIST argument contains one or more characters selected form a to f as well as A, each of which indicates the presence of an inner (endcyclic) double bond on the corresponding position.

     BONDLIST =
           none       :  cyclohexane
           a          :  1,2-double bond
           b          :  2,3-double bond
           c          :  4,3-double bond
           d          :  4,5-double bond
           e          :  5,6-double bond
           f          :  6,1-double bond
           A          :  aromatic circle

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

     SUBSLIST: list of substituents

for n = 1 to 6
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specifed) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)
                for the other ring fusion
nGB        :  beta single bond at n-atom (boldface)
                for the other ring fusion

e.g.
\cyclohexaneh{1==Cl;2==F}
\cyclohexaneh[c]{1==Cl;4==F;2==CH$_{3}$}
\cyclohexaneh[eb]{1D==O;4SA==MeO;4SB==OMe;5W==Cl;6==Cl}

\cyclohexaneh
21 \def\cyclohexaneh{\@ifnextchar[{\@cyclohexaneh[@}{\@cyclohexaneh[@]}}
22 \def\@cyclohexaneh[#1]#2{\sixheteroh[#1]{}{#2}}

The following definition has been deleted.

(Old defition)
% \begin{macro}{\cyclohexaneh}
%    \begin{macrocode}
\def\cyclohexaneh{\@ifnextchar[{\@cyclohexaneh[@}{\@cyclohexaneh[r]}}
\def\@cyclohexaneh[#1]#2{%
\iniatom\iniflag%initialize
\@reset@ylsw%
\ylpositionh{#2}{0}{0}{7}{0}%
\if@ylsw
\yl@shiftii=\@ylii
\yl@shifti=\@yli
\advance\yl@shiftii\yl@xdiff

 \advance\yl@shifti\yl@ydiff
 \begin{picture}(0,0)(-\yl@shiftii,-\yl@shifti)
 \reset@yl@xydiff%1999/1/6 by S. Fujita
\else
 \begin{picture}(880,800)(-\shifti,-\shiftii)
  \iforigpt \put(-\shifti,-\shiftii){\circle*{50}}%
           \put(-\noshift,-\noshift){\circle{50}}%
   \typeout{command `cyclohexaneh' origin: %
    (\the\noshift,\the\noshift) ---> (\the\shifti,\the\shiftii)}\fi%
\fi
  \put(0,0){\line(3,5){103}}%           %bond 1-2
  \put(0,0){\line(3,-5){103}}%          %bond 1-6
  \put(406,0){\line(-3,5){103}}%        %bond 4-3
  \put(406,0){\line(-3,-5){103}}%       %bond 4-3
  \put(103,171){\line(1,0){200}}%       %bond 2-3
  \put(103,-171){\line(1,0){200}}%      %bond 6-5
\@tfor\member:=#1\do{%
\if\member r%no endcyclic double bonds
\else\if\member a%
  \put(42,6){\line(3,5){78}}%           %double bond 1-2
\else\if\member b%
  \put(129,138){\line(1,0){148}}%       %double bond 2-3
\else\if\member c%
  \put(364,6){\line(-3,5){78}}%         %double bond 4-3
\else\if\member d%
  \put(364,-6){\line(-3,-5){78}}%       %double bond 4-5
\else\if\member e%
  \put(129,-138){\line(1,0){148}}%      %double bond 6-5
\else\if\member f%
  \put(42,-6){\line(3,-5){78}}%         %double bond 1-6
\else\if\member A%left aromatic circle
  \put(203,0){\circle{240}}%            %circle
\else
   \expandafter\twoCH@@R\member//%
   \set@fusionadd@sixh
\fi\fi\fi\fi\fi\fi\fi\fi}%
%%
\setsixringh{#2}{0}{0}{7}{0}%
\end{picture}}%               %end of \cyclohexaneh macro
%    \end{macrocode}
% \end{macro}

5 Benzene and benzoquinone derivatives

5.1 Vertical type

The macro \bzdrv is used for drawing benzene derivatives of vertical type as well as benzoquinone derivatives of vertical type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \cyclohexanev defined above.

The macro \bzdrv has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

 ****************************************
 * benzene and benzoquinone derivatives *
 *  (vertical type)                     *
 ****************************************
   \bzdrv[OPT]{SUBSLIST}

The OPT argument contains one character (r, l or c) for designate aromatic bond patterns or two characters (p? or o?) for quinone bond patterns.

     OPT: (bond pattern)
           none  or r :  right-handed set of double bonds
           l          :  left-handed set of double bonds
           c          :  aromatic circle
           p or pa    :  p-quinone (A)
           pb         :  p-quinone (B)
           pc         :  p-quinone (C)
           o or oa    :  o-quinone (A)
           ob         :  o-quinone (B)
           oc         :  o-quinone (C)
           od         :  o-quinone (D)
           oe         :  o-quinone (E)
           of         :  o-quinone (F)

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

     SUBSLIST: list of substituents
       for n = 1 to 6
           nD         :  exocyclic double bond at n-atom
           n or nS    :  exocyclic single bond at n-atom
           nA         :  alpha single bond at n-atom
           nB         :  beta single bond at n-atom
           nSA        :  alpha single bond at n-atom (boldface)

           nSB        :  beta single bond at n-atom (dotted line)
           nSa        :  alpha (not specified) single bond at n-atom
           nSb        :  beta (not specified) single bond at n-atom
           nSd        :  alpha single bond at n-atom (dotted line)
                           with an alternative direction to nSA
           nSu        :  beta single bond at n-atom (boldface)
                           with an alternative direction to nSB
           nFA        :  alpha single bond at n-atom (dotted line)
                           for ring fusion
           nFB        :  beta single bond at n-atom (boldface)
                           for ring fusion
           nGA        :  alpha single bond at n-atom (dotted line)
                           for the other ring fusion
           nGB        :  beta single bond at n-atom (boldface)
                           for the other ring fusion

       e.g.
        \bzdrv{1==Cl;2==F}
        \bzdrv[c]{1==Cl;4==F;2==CH$_{3}$}
        \bzdrv[pa]{1D==O;4D==O;2==CH$_{3}$}

\bzdrv
23 \def\bzdrv{\@ifnextchar[{\@bzdrv}{\@bzdrv[r]}}
24 \def\@bzdrv[#1]#2{%
25 \iforigpt \typeout{command `bzdrv' is based on `cyclohexanev'.}\fi%
26 %\expandafter\twoch@r#1{}%
27 \expandafter\twoCH@R#1//%
28 \if\@tmpa r% right-handed set of double bonds
29 \ifx\@tmpb\empty
30 \cyclohexanev[bdf]{#2}%
31 \fi
32 \else\if\@tmpa l%left-handed set of double bonds
33 \ifx\@tmpb\empty
34 \cyclohexanev[ace]{#2}%
35 \fi
36 \else\if\@tmpa c%aromatic circle
37 \ifx\@tmpb\empty
38 \cyclohexanev[A]{#2}%
39 \fi
40 \else\if\@tmpa p%p-quinone
41 \if\@tmpb a% (A)
42 \cyclohexanev[be]{#2}%
43 \else\ifx\@tmpb\empty% (A)
44 \cyclohexanev[be]{#2}%
45 \else\if\@tmpb b% (B)
46 \cyclohexanev[cf]{#2}%
47 \else\if\@tmpb c% (C)

48 \cyclohexanev[ad]{#2}%
49 \fi\fi\fi\fi%
50 \else\if\@tmpa o%o-quinone
51 \if\@tmpb a% (A)%
52 \cyclohexanev[ce]{#2}%
53 \else\ifx\@tmpb\empty% (A)
54 \cyclohexanev[ce]{#2}%
55 \else\if\@tmpb b%o-quinone (B)
56 \cyclohexanev[df]{#2}%
57 \else\if\@tmpb c%o-quinone (C)
58 \cyclohexanev[ae]{#2}%
59 \else\if\@tmpb d%o-quinone (D)
60 \cyclohexanev[bf]{#2}%
61 \else\if\@tmpb e%o-quinone (E)
62 \cyclohexanev[ac]{#2}%
63 \else\if\@tmpb f%o-quinone (F)
64 \cyclohexanev[bd]{#2}%
65 \fi\fi\fi\fi\fi\fi\fi%
66 \else% added 1998/11/15 by S. Fujita
67 \cyclohexanev[@#1]{#2}%
68 \fi\fi\fi\fi\fi%
69 }%               %end of \bzdrv macro

5.2 Horizontal type

The macro \bzdrh is used for drawing benzene derivatives of vertical type as well as benzoquinone derivatives of vertical type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \cyclohexanev defined above.

The macro \bzdrh has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

 ****************************************
 * benzene and benzoquinone derivatives *
 *  (horizontal type)                   *
 ****************************************
   \bzdrh[OPT]{SUBSLIST}

The OPT argument contains one character (r, l or c) for designate aromatic bond patterns or two characters (p? or o?) for quinone bond patterns.

     OPT: (bond pattern)
           none  or r :  right-handed set of double bonds
           l          :  left-handed set of double bonds
           c          :  aromatic circle

p or pa    :  p-quinone (A)
pb         :  p-quinone (B)
pc         :  p-quinone (C)
o or oa    :  o-quinone (A)
ob         :  o-quinone (B)
oc         :  o-quinone (C)
od         :  o-quinone (D)
oe         :  o-quinone (E)
of         :  o-quinone (F)

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 6
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)
                for the other ring fusion
nGB        :  beta single bond at n-atom (boldface)
                for the other ring fusion

e.g.
\bzdrh{1==Cl;2==F}
\bzdrh[c]{1==Cl;4==F;2==CH$_{3}$}

        \bzdrh[pa]{1D==O;4D==O;2==CH$_{3}$}

\bzdrh
70 \def\bzdrh{\@ifnextchar[{\@bzdrh}{\@bzdrh[r]}}
71 \def\@bzdrh[#1]#2{%
72 \iforigpt \typeout{command `bzdrh' is based on `cyclohexaneh'.}\fi%
73 %\expandafter\twoch@r#1{}%
74 \expandafter\twoCH@R#1//%
75 %\begin{picture}(800,880)(-\noshift,-\noshift)
76 \if\@tmpa r% right-handed set of double bonds
77 \cyclohexaneh[bdf]{#2}%
78 \else\if\@tmpa l%left-handed set of double bonds
79 \cyclohexaneh[ace]{#2}%
80 \else\if\@tmpa c%aromatic circle
81 \cyclohexaneh[A]{#2}%
82 \else\if\@tmpa p%p-quinone
83 \if\@tmpb a% (A)
84 \cyclohexaneh[be]{#2}%
85 \else\ifx\@tmpb\empty% (A)
86 \cyclohexaneh[be]{#2}%
87 \else\if\@tmpb b% (B)
88 \cyclohexaneh[cf]{#2}%
89 \else\if\@tmpb c% (C)
90 \cyclohexaneh[ad]{#2}%
91 \fi\fi\fi\fi%
92 \else\if\@tmpa o%o-quinone
93 \if\@tmpb a% (A)%
94 \cyclohexaneh[ce]{#2}%
95 \else\ifx\@tmpb\empty% (A)
96 \cyclohexaneh[ce]{#2}%
97 \else\if\@tmpb b%o-quinone (B)
98 \cyclohexaneh[df]{#2}%
99 \else\if\@tmpb c%o-quinone (C)
100 \cyclohexaneh[ae]{#2}%
101 \else\if\@tmpb d%o-quinone (D)
102 \cyclohexaneh[bf]{#2}%
103 \else\if\@tmpb e%o-quinone (E)
104 \cyclohexaneh[ac]{#2}%
105 \else\if\@tmpb f%o-quinone (F)
106 \cyclohexaneh[bd]{#2}%
107 \fi\fi\fi\fi\fi\fi\fi%
108 \else% added 1998/11/15 by S. Fujita
109 \cyclohexaneh[@#1]{#2}%
110 \fi\fi\fi\fi\fi%
111 %\end{picture}
112 }%               %end of \bzdrh macro

6 Decaline derivatives

6.1 Vertical type

The macro \decalinev is used for drawing decaline derivatives of vertical type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \setdecaringv.

 ************************
 * decaline derivatives *
 *  (vertical type)     *
 ************************
 The following numbering is adopted in this macro.
          8 (0G)1
          *  8a *
     7  *   * *   *  2
       |     |     |
       |     |     |
     6  *   * *   *  3
          *  4a *
          5 (0F)4
          ^
          |
          the original point

The macro \decalinev has an argument SUBSLIST as well as an optional argument BONDLIST.

   \decalinev[BONDLIST]{SUBSLIST}

The BONDLIST argument contains one or more characters selected form a to k, each of which indicates the presence of an inner (endcyclic) double bond on the corresponding position. The character A or B indicates aromatic bond patterns (circles).

     BONDLIST =
           none       :  decaline
           a          :  1,2-double bond
           b          :  2,3-double bond
           c          :  3,4-double bond
           d          :  4,4a-double bond
           e          :  5,4a-double bond
           f          :  5,6-double bond
           g          :  6,7-double bond
           h          :  8,7-double bond
           i          :  8,8a-double bond

j          :  1,8a-double bond
k          :  4a,8a-double bond
K          :  4a,8a-double bond in the other ring
A          :  left aromatic circle
B          :  right aromatic circle

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 8
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom

nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)
                for the other ring fusion
nGB        :  beta single bond at n-atom (boldface)
                for the other ring fusion
for 0 (fused positions)
0FA        :  alpha single bond at 8a
0FB        :  beta single bond at 8a
0FU        :  unspecified single bond at 8a
0GA        :  alpha single bond at 4a
0GB        :  beta single bond at 4a
0GU        :  unspecified single bond at 4a

      e.g.
       \decalinev{1==Cl;2==F;0FA==;0GA==}
       \decalinev[c]{1==Cl;4==F;2==CH$_{3}$}
       \decalinev[eb]{1D==O;4SA==MeO;4SB==OMe;5W==Cl;6==Cl;7==Cl;8==Cl}

The macro \decalinev has been changed into a subsidiary macro of \decaheterov (version 2.00).

\decalinev
113 \def\decalinev{\@ifnextchar[{\@decalinev[@}{\@decalinev[@]}}
114 \def\@decalinev[#1]#2{%
115 \iforigpt \typeout{command `decalinev' is based on `decaheterov'.}\fi%
116 \decaheterov[#1]{}{#2}}

The following definition has been stopped using.

(Old definition)
% \begin{macro}{\decalinev}
%    \begin{macrocode}
\def\decalinev{\@ifnextchar[{\@decalinev[@}{\@decalinev[r]}}
\def\@decalinev[#1]#2{%
\iniflag\iniatom
\@reset@ylsw\reset@@yl
\ylposition{#2}{-7}{0}{2}{0}% for 8
\if@ylsw \ifx\@@ylii\empty
\def\@@ylii{0}\def\@@yli{0}\fi
\else\ylposition{#2}{-1}{3}{7}{0}% for 5 to 7
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{0}\def\@@yli{0}\fi\else
\ylposition{#2}{0}{0}{5}{0}% for 1 to 4
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{-342}\def\@@yli{0}\fi\else
\ylfusedposition{#2}{0}{-1}{1}{0}% for 4a and 8a
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{0}\def\@@yli{0}\fi\fi
\if@ylsw
 \yl@shiftii=\@ylii
 \ifx\@@ylii\empty\else
 \advance\yl@shiftii\@@ylii\fi
 \yl@shifti=\@yli
 \ifx\@@yli\empty\else
 \advance\yl@shifti\@@yli\fi
 \advance\yl@shiftii\yl@xdiff

\advance\yl@shifti\yl@ydiff
\begin{picture}(0,0)(-\yl@shiftii,-\yl@shifti)
\reset@yl@xydiff%1999/1/6 by S. Fujita
\else
\begin{picture}(1200,880)(-\shiftii,-\shifti)
\iforigpt \put(-\shiftii,-\shifti){\circle*{50}}%
         \put(-\noshift,-\noshift){\circle{50}}%
 \typeout{command `decalinev' origin: %
  (\the\noshift,\the\noshift) ---> (\the\shiftii,\the\shifti)}\fi%
\fi
\put(0,406){\line(-5,-3){171}}%       %bond 8-7     1-6
\put(0,406){\line(5,-3){171}}%        %bond 8-8a    1-2
\put(0,0){\line(-5,3){171}}%          %bond 5-6     4-5
\put(0,0){\line(5,3){171}}%           %bond 5-4a    4-3
\put(171,103){\line(0,1){200}}%       %bond 4a-8a   3-2
\put(-171,103){\line(0,1){200}}%      %bond 5-7     5-6
\put(342,406){\line(-5,-3){171}}%     %bond 1-8a    1-6
\put(342,406){\line(5,-3){171}}%      %bond 1-2     1-2
\put(342,0){\line(-5,3){171}}%        %bond 4-4a    4-5
\put(342,0){\line(5,3){171}}%         %bond 4-3     4-3
\put(513,103){\line(0,1){200}}%       %bond 3-2     3-2
\@tfor\member:=#1\do{%
\if\member r%no endcyclic double bonds
\else\if\member a%
\put(348,364){\line(5,-3){126}}%      %double bond 1-2
\else\if\member b%
\put(482,129){\line(0,1){148}}%       %double bond 3-2
\else\if\member c%
\put(348,42){\line(5,3){126}}%        %double bond 4-3
\else\if\member d%
\put(348,42){\line(-5,3){126}}%       %double bond 4-4a
\else\if\member e%
\put(6,42){\line(5,3){126}}%          %double bond 5-4a
\else\if\member f%
\put(-6,42){\line(-5,3){126}}%        %double bond 5-6
\else\if\member g%
\put(-138,129){\line(0,1){148}}%      %double bond 6-7
\else\if\member h%
\put(-6,364){\line(-5,-3){126}}%      %double bond 8-7
\else\if\member i%
\put(6,364){\line(5,-3){126}}%        %double bond 8-8a
\else\if\member j%
\put(348,364){\line(-5,-3){126}}%     %double bond 1-8a
\else\if\member k%
\put(138,129){\line(0,1){148}}%       %double bond 4a-8a
\else\if\member K%
\put(204,129){\line(0,1){148}}%       %double bond 4a-8a in the other ring
\else\if\member A%left aromatic circle
\put(0,203){\circle{240}}%            %circle
\else\if\member B%right aromatic circle

  \put(342,203){\circle{240}}%          %circle
\else
% ***********************
% * setting fused rings *
% ***********************
   \expandafter\twoCH@R\member//%
   \set@fusion@decav
\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi}%
%    \end{macrocode}
%
% The command |\setdecaringv{#2}{0}{-1}{9}{0}|
% has been stopped to use and replaced
% by the following set of commands.
%
%    \begin{macrocode}
\put(0,0){\setsixringv{#2}{-7}{0}{2}{0}}% for 8
\put(0,0){\setsixringv{#2}{-1}{3}{7}{0}}% for 5 to 7
\put(0,0){\setsixringv{#2}{-1}{3}{7}{0}}% for 5 to 7
\put(342,0){\setsixringv{#2}{0}{0}{5}{0}}% for 1 to 4
\put(0,0){\setfusedbond{#2}{0}{-1}{1}{0}}% for 4a and 8a
\end{picture}}%               %end of \decalinev macro
%    \end{macrocode}
% \end{macro}

6.2 Vertical-bottom type

The macro \decalinevb is used for drawing decaline derivatives of vertical-bottom type (numbering from the bottom to the left-upper part). The word “vertical” means that each benzene ring is a vertical type. The word “bottom” means that the benzene ring with young locant numbers is located at the bottom. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \setsixringv.

***************************
* decaline derivatives    *
*  (vertical bottom type) *
***************************
The following numbering is adopted in this macro.
         7
         *
    6  *   * 8
      |     |
      |     | 0G (4a)
    5  *   * *
 0F(4a)  *     * 1
         |     |
         |     |

        4  *   *  2
             *
             3
             ^
             |
            the original point

The macro \decalinevb has an argument SUBSLIST as well as an optional argument BONDLIST.

   \decalinevb[BONDLIST]{SUBSLIST}

The BONDLIST argument contains one or more characters selected form a to k, each of which indicates the presence of an inner (endcyclic) double bond on the corresponding position. The character A or B indicates aromatic bond patterns (circles).

     BONDLIST =
           none       :  decaline
           a          :  1,2-double bond
           b          :  2,3-double bond
           c          :  3,4-double bond
           d          :  4,4a-double bond
           e          :  5,4a-double bond
           f          :  5,6-double bond
           g          :  6,7-double bond
           h          :  8,7-double bond
           i          :  8,8a-double bond
           j          :  1,8a-double bond
           k          :  4a,8a-double bond
           K          :  4a,8a-double bond in the other ring
           A          :  left aromatic circle
           B          :  right aromatic circle

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

     SUBSLIST: list of substituents
       for n = 1 to 8
           nD         :  exocyclic double bond at n-atom
           n or nS    :  exocyclic single bond at n-atom

nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)
                for the other ring fusion
nGB        :  beta single bond at n-atom (boldface)
                for the other ring fusion
for 0 (fused positions)
0FA        :  alpha single bond at 8a
0FB        :  beta single bond at 8a
0FU        :  unspecified single bond at 8a
0GA        :  alpha single bond at 4a
0GB        :  beta single bond at 4a
0GU        :  unspecified single bond at 4a

e.g.
\decalinevb{1==Cl;2==F;0FA==;0GA==}
\decalinevb[c]{1==Cl;4==F;2==CH$_{3}$}
\decalinevb[eb]{1D==O;4SA==MeO;4SB==OMe;5W==Cl;6==Cl;7==Cl;8==Cl}

The macro \decalinevb has been changed into a subsidiary macro of \decaheterovb (version 2.00).

\decalinevb
117 \def\decalinevb{\@ifnextchar[{\@decalinevb[@}{\@decalinevb[@]}}
118 \def\@decalinevb[#1]#2{%
119 \iforigpt \typeout{command `decalinevb' is based on `decaheterovb'.}\fi%
120 \decaheterovb[#1]{}{#2}}

The following old definition has been stopped using.

(Old definition)
% \begin{macro}{\decalinevb}
%    \begin{macrocode}
\def\decalinevb{\@ifnextchar[{\@decalinevb[@}{\@decalinevb[r]}}
\def\@decalinevb[#1]#2{%
\iniflag\iniatom
\@reset@ylsw\reset@@yl
\ylposition{#2}{1}{1}{6}{0}% for 1 to 4
\if@ylsw \ifx\@@ylii\empty
\def\@@ylii{0}\def\@@yli{0}\fi
\else\ylposition{#2}{0}{4}{7}{0}% for 5 to 6
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{171}\def\@@yli{-303}\fi\else
\ylposition{#2}{-6}{0}{3}{0}% for 7 to 8
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{171}\def\@@yli{-303}\fi\else
\ylfusedposition{#2}{2}{1}{3}{-1}% for 4a and 8a
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{0}\def\@@yli{0}\fi\fi
\if@ylsw
 \yl@shiftii=\@ylii
 \ifx\@@ylii\empty\else
 \advance\yl@shiftii\@@ylii\fi
 \yl@shifti=\@yli
 \ifx\@@yli\empty\else
 \advance\yl@shifti\@@yli\fi
 \advance\yl@shiftii\yl@xdiff
 \advance\yl@shifti\yl@ydiff
 \begin{picture}(0,0)(-\yl@shiftii,-\yl@shifti)
 \reset@yl@xydiff%1999/1/6 by S. Fujita
\else
 \begin{picture}(1000,1200)(-550,-240)
  \iforigpt \put(-550,-240){\circle*{50}}%
           \put(0,0){\circle{50}}%
   \typeout{command `decalinevb' origin: %
    (0,0) ---> (550,240)}\fi%
\fi
  \put(0,0){\line(5,3){171}}%           %bond 2-3
  \put(0,0){\line(-5,3){171}}%          %bond 4-3
  \put(171,103){\line(0,1){200}}%       %bond 2-1
  \put(-171,103){\line(0,1){200}}%      %bond 4-4a
  \put(0,406){\line(-5,-3){171}}%       %bond 8a-1
  \put(0,406){\line(5,-3){171}}%        %bond 8a-4a
  \put(0,406){\line(0,1){200}}%         %bond 8a-8
  \put(-342,406){\line(0,1){200}}%      %bond 5-6     3-2
  \put(-171,303){\line(-5,3){171}}%     %bond 4a-5
  \put(-171,709){\line(-5,-3){171}}%    %bond 7-6

  \put(-171,709){\line(5,-3){171}}%     %bond 7-8
\@tfor\member:=#1\do{%
\if\member r%no endcyclic double bonds
\else\if\member a%
  \put(140,129){\line(0,1){148}}%       %double bond 1-2
\else\if\member b%
  \put(6,42){\line(5,3){126}}%          %double bond 2-3
\else\if\member c%
  \put(-6,42){\line(-5,3){126}}%        %double bond 3-4
\else\if\member d%
  \put(-140,129){\line(0,1){148}}%      %double bond 4-4a
\else\if\member e%
  \put(-177,345){\line(-5,3){126}}%     %double bond 4a-5
\else\if\member f%
  \put(-311,432){\line(0,1){148}}%      %double bond 5-6
\else\if\member g%
  \put(-177,667){\line(-5,-3){126}}%    %double bond 7-6
\else\if\member h%
  \put(-165,667){\line(5,-3){126}}%     %double bond 7-8
\else\if\member i%
  \put(-31,432){\line(0,1){148}}%       %double bond 8a-8
\else\if\member j%
  \put(6,364){\line(5,-3){126}}%        %double bond 8a-1
\else\if\member k%
  \put(-6,364){\line(-5,-3){126}}%      %double bond 8a-4a
\else\if\member K%
  \put(-165,345){\line(5,3){126}}%     %double bond 8a-4a on the other ring
\else\if\member A%left aromatic circle
  \put(-171,506){\circle{240}}%         %circle
\else\if\member B%right aromatic circle
  \put(0,203){\circle{240}}%            %circle
\else
% ***********************
% * setting fused rings *
% ***********************
   \expandafter\twoCH@R\member//%
   \set@fusion@decavb
\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi}%
% ************************
% * setting substituents *
% ************************
\put(0,0){\setsixringv{#2}{1}{1}{6}{0}}% for 1 to 4
\put(-171,303){\setsixringv{#2}{0}{4}{7}{0}}% for 5 to 6
\put(-171,303){\setsixringv{#2}{-6}{0}{3}{0}}% for 7 to 8
\put(0,0){\setfusedbond{#2}{2}{1}{3}{-1}}% for 4a and 8a
\end{picture}}%               %end of \decalinevb macro
%    \end{macrocode}
% \end{macro}

6.3 Vertical-top type

The macro \decalinevt is used for drawing decaline derivatives of vertical-bottom type (numbering from the top to the left-down part). The word “vertical” means that each benzene ring is a vertical type. The word “top” means that the benzene ring with young locant numbers is located at the top. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \setsixringv.

 ************************
 * decaline derivatives *
 *  (vertical-top type) *
 ************************
 The following numbering is adopted in this macro.
                     2
                     *
                1  *   * 3
                  |     |
                  |     |
         0G (8a)   *   * 4
             8   *   * 0F(4a)
                |     |
                |     |
              7  *   *  5
                   *
                   6
                   ^
                   |
                   the original point

The macro \decalinevt has an argument SUBSLIST as well as an optional argument BONDLIST.

   \decalinevt[BONDLIST]{SUBSLIST}

The BONDLIST argument contains one or more characters selected form a to k, each of which indicates the presence of an inner (endcyclic) double bond on the corresponding position. The character A or B indicates aromatic bond patterns (circles).

     BONDLIST =
           none       :  decaline
           a          :  1,2-double bond
           b          :  2,3-double bond
           c          :  3,4-double bond
           d          :  4,4a-double bond

e          :  5,4a-double bond
f          :  5,6-double bond
g          :  6,7-double bond
h          :  8,7-double bond
i          :  8,8a-double bond
j          :  1,8a-double bond
k          :  4a,8a-double bond
K          :  4a,8a-double bond in the other ring
A          :  left aromatic circle
B          :  right aromatic circle

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 8
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)
                for the other ring fusion
nGB        :  beta single bond at n-atom (boldface)
                for the other ring fusion
for 0 (fused positions)
0FA        :  alpha single bond at 8a
0FB        :  beta single bond at 8a
0FU        :  unspecified single bond at 8a
0GA        :  alpha single bond at 4a

          0GB        :  beta single bond at 4a
          0GU        :  unspecified single bond at 4a

      e.g.
       \decalinevt{1==Cl;2==F;0FA==;0GA==}
       \decalinevt[c]{1==Cl;4==F;2==CH$_{3}$}
       \decalinevt[eb]{1D==O;4SA==MeO;4SB==OMe;5W==Cl;6==Cl;7==Cl;8==Cl}

The macro \decalinevt has been changed into a subsidiary macro of \decaheterovt (version 2.00).

\decalinevt
121 \def\decalinevt{\@ifnextchar[{\@decalinevt[@}{\@decalinevt[@]}}
122 \def\@decalinevt[#1]#2{%
123 \iforigpt \typeout{command `decalinevt' is based on `decaheterovt'.}\fi%
124 \decaheterovt[#1]{}{#2}}

The following old definition has been stopped using.

(Old definition)
% \begin{macro}{\decalinevt}
%    \begin{macrocode}
\def\decalinevt{\@ifnextchar[{\@decalinevt[@}{\@decalinevt[r]}}
\def\@decalinevt[#1]#2{%
\iniflag\iniatom
\@reset@ylsw\reset@@yl
\ylposition{#2}{5}{5}{7}{0}% for 1
\if@ylsw \ifx\@@ylii\empty
\def\@@ylii{-171}\def\@@yli{-303}\fi
\else\ylposition{#2}{-1}{0}{4}{0}% for 2 to 4
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{-171}\def\@@yli{-303}\fi\else
\ylposition{#2}{-2}{2}{7}{0}% for 5 to 8
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{0}\def\@@yli{0}\fi\else
\ylfusedposition{#2}{1}{0}{2}{-1}% for 4a and 8a
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{0}\def\@@yli{0}\fi\fi
\if@ylsw
 \yl@shiftii=\@ylii
 \ifx\@@ylii\empty\else
 \advance\yl@shiftii\@@ylii\fi

\yl@shifti=\@yli
\ifx\@@yli\empty\else
\advance\yl@shifti\@@yli\fi
\advance\yl@shiftii\yl@xdiff
\advance\yl@shifti\yl@ydiff
\begin{picture}(0,0)(-\yl@shiftii,-\yl@shifti)
\reset@yl@xydiff%1999/1/6 by S. Fujita
\else
\begin{picture}(1000,1200)(-400,-240)
\iforigpt \put(-400,-240){\circle*{50}}%
         \put(0,0){\circle{50}}%
 \typeout{command `decalinevb' origin: %
  (0,0) ---> (400,240)}\fi%
\fi
\put(0,0){\line(5,3){171}}%           %bond 6-5
\put(0,0){\line(-5,3){171}}%          %bond 6-7
\put(171,103){\line(0,1){200}}%       %bond 5-4a
\put(-171,103){\line(0,1){200}}%      %bond 7-8
\put(0,406){\line(-5,-3){171}}%       %bond 8a-8
\put(0,406){\line(5,-3){171}}%        %bond 8a-4a
\put(0,406){\line(0,1){200}}%         %bond 8a-1
\put(342,406){\line(0,1){200}}%       %bond 4-3
\put(171,303){\line(5,3){171}}%       %bond 4a-4
\put(171,709){\line(-5,-3){171}}%     %bond 2-1
\put(171,709){\line(5,-3){171}}%      %bond 2-3
\@tfor\member:=#1\do{%
\if\member r%no endcyclic double bonds
\else\if\member a%
\put(165,667){\line(-5,-3){126}}%    %double bond 2-1
\else\if\member b%
\put(177,667){\line(5,-3){126}}%     %double bond 2-3
\else\if\member c%
\put(311,432){\line(0,1){148}}%      %double bond 4-3
\else\if\member d%
\put(177,345){\line(5,3){126}}%      %double bond 4a-4
\else\if\member e%
\put(140,129){\line(0,1){148}}%      %double bond 5-4a
\else\if\member f%
\put(6,42){\line(5,3){126}}%         %double bond 6-5
\else\if\member g%
\put(-6,42){\line(-5,3){126}}%       %double bond 6-7
\else\if\member h%
\put(-140,129){\line(0,1){148}}%     %double bond 7-8
\else\if\member i%
\put(-6,364){\line(-5,-3){126}}%     %double bond 8a-8
\else\if\member j%
\put(31,432){\line(0,1){148}}%       %double bond 8a-1
\else\if\member k%
\put(6,364){\line(5,-3){126}}%       %double bond 8a-4a
\else\if\member K%

  \put(165,345){\line(-5,3){126}}%     %double bond 8a-4a on the other ring
\else\if\member A%left aromatic circle
  \put(171,506){\circle{240}}%         %circle
\else\if\member B%right aromatic circle
  \put(0,203){\circle{240}}%            %circle
\else
% ***********************
% * setting fused rings *
% ***********************
  \expandafter\twoCH@R\member//%
   \set@fusion@decavt
\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi}%
% %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
% % setting substituents and exocyclic bonds %
% %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\put(171,303){\setsixringv{#2}{5}{5}{7}{0}}% for 1
\put(171,303){\setsixringv{#2}{-1}{0}{4}{0}}% for 2 to 4
\put(0,0){\setsixringv{#2}{-2}{2}{7}{0}}% for 5 to 8
\put(0,0){\setfusedbond{#2}{1}{0}{2}{-1}}% for 4a and 8a
\end{picture}}%               %end of \decalinevt macro
%    \end{macrocode}
% \end{macro}

6.4 Horizontal type

The macro \decalineh is used for drawing decaline derivatives of vertical type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macros \setsixringh and

\setfusedbond.
************************
* decaline derivatives *
*  (horizontal type)   *
************************
The following numbering is adopted in this macro.
                               2      3
                                -----
                              *       *
                          1  *          * 4
                              *       *
                       8a (0G)  -----  4a (0F)
                              *       *
  the original point ===> 8 *           * 5
         (0,0)                *       *
                                -----
                               7      6

The macro \decalineh has an argument SUBSLIST as well as an optional argument BONDLIST.

\decalineh[BONDLIST]{SUBSLIST}

The BONDLIST argument contains one or more characters selected form a to k, each of which indicates the presence of an inner (endcyclic) double bond on the corresponding position. The character A or B indicates aromatic bond patterns (circles).

BONDLIST =
none       :  decaline
a          :  1,2-double bond
b          :  2,3-double bond
c          :  3,4-double bond
d          :  4,4a-double bond
e          :  5,4a-double bond
f          :  5,6-double bond
g          :  6,7-double bond
h          :  8,7-double bond
i          :  8,8a-double bond
j          :  1,8a-double bond
k          :  4a,8a-double bond
K          :  4a,8a-double bond in the other ring
A          :  left aromatic circle
B          :  right aromatic circle

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 8
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom

nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)
                for the other ring fusion
nGB        :  beta single bond at n-atom (boldface)
                for the other ring fusion
for 0 (fused positions)
0FA        :  alpha single bond at 8a
0FB        :  beta single bond at 8a
0FU        :  unspecified single bond at 8a
0GA        :  alpha single bond at 4a
0GB        :  beta single bond at 4a
0GU        :  unspecified single bond at 4a

e.g.
\decalineh{1==Cl;2==F;0FA==;0GA==}
\decalineh[c]{1==Cl;4==F;2==CH$_{3}$}
\decalineh[eb]{1D==O;4SA==MeO;4SB==OMe;5W==Cl;6==Cl;7==Cl;8==Cl}

The macro \decalineh has been changed into a subsidiary macro of \decaheteroh (version 2.00).

\decalineh
125 \def\decalineh{\@ifnextchar[{\@decalineh[@}{\@decalineh[@]}}
126 \def\@decalineh[#1]#2{%
127 \iforigpt \typeout{command `decalineh' is based on `decaheteroh'.}\fi%
128 \decaheteroh[#1]{}{#2}}

The following old definition has been stopped using.

(Old defintion)
% \begin{macro}{\decalineh}
%    \begin{macrocode}
\def\decalineh{\@ifnextchar[{\@decalineh[@}{\@decalineh[r]}}
\def\@decalineh[#1]#2{%
\iniflag\iniatom
\@reset@ylsw\reset@@yl
\ylpositionh{#2}{-7}{0}{2}{0}% for 8
\if@ylsw \ifx\@@ylii\empty

\def\@@ylii{0}\def\@@yli{0}\fi
\else\ylpositionh{#2}{-1}{3}{7}{0}% for 5 to 7
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{0}\def\@@yli{0}\fi\else
\ylpositionh{#2}{0}{0}{5}{0}% for 1 to 4
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{0}\def\@@yli{-342}\fi\else
\ylfusedposition{#2}{3}{2}{4}{0}% for 4a and 8a
\fi
\if@ylsw\ifx\@@ylii\empty
\def\@@ylii{0}\def\@@yli{0}\fi\fi
\if@ylsw
 \yl@shiftii=\@ylii
 \ifx\@@ylii\empty\else
 \advance\yl@shiftii\@@ylii\fi
 \yl@shifti=\@yli
 \ifx\@@yli\empty\else
 \advance\yl@shifti\@@yli\fi
 \advance\yl@shiftii\yl@xdiff
 \advance\yl@shifti\yl@ydiff
 \begin{picture}(0,0)(-\yl@shiftii,-\yl@shifti)
 \reset@yl@xydiff%1999/1/6 by S. Fujita
\else
 \begin{picture}(880,1200)(-\shifti,-\shiftii)
  \iforigpt \put(-\shifti,-\shiftii){\circle*{50}}%
           \put(\noshift,\noshift){\circle{50}}%
   \typeout{command `decalineh' origin: %
    (\the\noshift,\the\noshift) ---> (\the\shifti,\the\shiftii)}\fi%
\fi
  \put(0,0){\line(3,5){103}}%            %bond 8-8a
  \put(0,0){\line(3,-5){103}}%          %bond 8-7
  \put(406,0){\line(-3,5){103}}%        %bond 4a-5
  \put(406,0){\line(-3,-5){103}}%       %bond 5-6
  \put(103,171){\line(1,0){200}}%       %bond 4a-8a
  \put(103,-171){\line(1,0){200}}%      %bond 7-8
  \put(0,342){\line(3,5){103}}%         %bond 1-2
  \put(0,342){\line(3,-5){103}}%        %bond 1-6
  \put(406,342){\line(-3,5){103}}%      %bond 4-3
  \put(406,342){\line(-3,-5){103}}%     %bond 4-4a
  \put(103,513){\line(1,0){200}}%       %bond 2-3
\@tfor\member:=#1\do{%
\if\member r%no endcyclic double bonds
\else\if\member a%
  \put(42,348){\line(3,5){78}}%         %double bond 1-2
\else\if\member b%
  \put(129,480){\line(1,0){148}}%       %double bond 2-3
\else\if\member c%
  \put(364,348){\line(-3,5){78}}%       %double bond 4-3

\else\if\member d%
  \put(364,348){\line(-3,-5){78}}%      %double bond 4-4a
\else\if\member e%
  \put(364,6){\line(-3,5){78}}%         %double bond 4a-5
\else\if\member f%
  \put(364,-6){\line(-3,-5){78}}%       %double bond 5-6
\else\if\member g%
  \put(129,-138){\line(1,0){148}}%      %double bond 6-7
\else\if\member h%
  \put(42,-6){\line(3,-5){78}}%         %double bond 7-8
\else\if\member i%
  \put(42,6){\line(3,5){78}}%           %double bond 8-9a
\else\if\member j%
  \put(42,348){\line(3,-5){78}}%        %double bond 1-8a
\else\if\member k%
  \put(129,138){\line(1,0){148}}%       %double bond 8a-4a
\else\if\member K%
  \put(129,204){\line(1,0){148}}%       %double bond 8a-4a
\else\if\member A%left aromatic circle
  \put(203,0){\circle{240}}%            %circle
\else\if\member B%right aromatic circle
  \put(203,342){\circle{240}}%          %circle
% ***********************
% * setting fused rings *
% ***********************
\else
   \expandafter\twoCH@R\member//%
   \put(0,342){\set@fusion@decah}%
\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi}%
%%
\put(0,0){\setsixringh{#2}{-7}{0}{2}{0}}% for 8
\put(0,0){\setsixringh{#2}{-1}{3}{7}{0}}% for 5 to 7
\put(0,342){\setsixringh{#2}{0}{0}{5}{0}}% for 1 to 4
\put(0,0){\setfusedbond{#2}{3}{2}{4}{0}}% for 4a and 8a
\end{picture}}%               %end of \decalineh macro
%    \end{macrocode}
% \end{macro}

7 Naphthalene and naphtoquinone derivatives

7.1 Vertical type

The macro \naphdrv is used for drawing naphthalene derivatives of vertical type as well as naphthoquinone derivatives of vertical type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \decalinev.

The macro \naphdrv has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

 **********************************************
 * naphthalene and naphthoquinone derivatives *
 *  (vertical type)                           *
 **********************************************
   \naphdrv[OPT]{SUBSLIST}

The OPT argument contains one or two characters.

The character ‘A’ is used for designating an aromatic bond pattern, in which each of the six-membered rings has a central circle.

     OPT (bond pattern)
           none       :  naphthalene
           A          :  aromatic circle

One character ‘p’ or two characters (p?) in the OPT argument designate a pquinone nucleus, which is present in the right-hand ring for the case of ‘p’ or ‘pa’ and in the left-hand ring for the case of ‘pb’.

           p or pa    :  1,4-quinone (A) left aromatic, right quinone
           pb         :  1,4-quinone (B) right aromatic, left quinone

One character ‘o’ or two characters (o?) in the OPT argument designate an o-quinone nucleus in either side of the fused ring. The second character of each argument indicates one of the possible type of bond patterns.

           o or oa    :  o-quinone (A)
           ob         :  o-quinone (B)
           oc         :  o-quinone (C)
           od         :  o-quinone (D)
           oe         :  o-quinone (E)
           of         :  o-quinone (F)

One character ‘q’ or two characters (q?) in the OPT argument designate an 2,6-, 1,5- or 1,7-quinone nucleus. The second character of each argument indicates one of the possible type of bond patterns.

           q or qa    :  2,6-quinone (A)
           qb         :  2,6-quinone (B)  3,7
           qc         :  1,5-quinone (C)
           qd         :  1,5-quinone (D)  4,8

qe         :  1,7-quinone (E)
qf         :  1,7-quinone (F)  2,8
qg         :  1,7-quinone (G)  4,6
qh         :  1,7-quinone (H)  3,5

One character ‘P’ or two characters (P?) in the OPT argument designate a 1,4,5,8- or 1,2,5,8-quinone nucleus.

P or Pa    :  1,4,5,8-quinone (A)
Pb         :  1,2,5,8-quinone (B)

One character ‘Q’ or ‘O’ or two characters (O?) in the OPT argument designate quinone nuclei of remaining types.

Q          :  1,2,3,4-quinone
O or Oa    :  1,2,5,6-quinone (A)
Ob         :  1,2,7,8-quinone (B)
Oc         :  1,2,3,5-quinone (C)
Od         :  1,2,3,7-quinone (D)

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 8
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion

            nFB        :  beta single bond at n-atom (boldface)
                            for ring fusion
            nGA        :  alpha single bond at n-atom (dotted line)
                            for the other ring fusion
            nGB        :  beta single bond at n-atom (boldface)
                            for the other ring fusion

        e.g.
         \naphdrv{1==Cl;2==F}
         \naphdrv[A]{1==Cl;4==F;2==CH$_{3}$}

\naphdrv
129 \def\naphdrv{\@ifnextchar[{\@naphdrv}{\@naphdrv[r]}}
130 \def\@naphdrv[#1]#2{%
131 \iforigpt \typeout{command `naphdrv' is based on `decalinev'.}\fi%
132 \expandafter\threech@r#1{}{}%
133 \if\@tmpa r% right-handed set of double bonds
134 \decalinev[acfhk]{#2}%
135 \else\if\@tmpa A%aromatic circle
136 \decalinev[AB]{#2}%

137 %

138 \else\if\@tmpa p%p-quinone
139 \if\@tmpb a% (A)
140 \decalinev[bfhk]{#2}%
141 \else\ifx\@tmpb\empty% (A)
142 \decalinev[bfhk]{#2}%
143 \else\if\@tmpb b% (B)
144 \decalinev[acgk]{#2}%
145 \fi\fi\fi%
146 \else%
147 \if\@tmpa o%o-quinone
148 \if\@tmpb a% (A)
149 \decalinev[cfhk]{#2}%
150 \else\ifx\@tmpb\empty% (A)
151 \decalinev[cfhk]{#2}%
152 \else\if\@tmpb b%o-quinone (B)
153 \decalinev[dfhj]{#2}%
154 \else\if\@tmpb c%o-quinone (C)
155 \decalinev[afhk]{#2}%
156 \else\if\@tmpb d%o-quinone (D)
157 \decalinev[achk]{#2}%
158 \else\if\@tmpb e%o-quinone (E)
159 \decalinev[acei]{#2}%
160 \else\if\@tmpb f%o-quinone (F)
161 \decalinev[acfk]{#2}%
162 \fi\fi\fi\fi\fi\fi%
163 \fi%
164 %%%

165 \else\if\@tmpa q%various quinone
166 \if\@tmpb a% (A)2,6-type
167 \decalinev[cehj]{#2}%
168 \else\ifx\@tmpb\empty% (A)2,6-type
169 \decalinev[cehj]{#2}%
170 \else\if\@tmpb b%2,6-quinone (B)
171 \decalinev[adfi]{#2}%
172 \else\if\@tmpb c%1,5-quinone (C)
173 \decalinev[bdgi]{#2}%
174 \else\if\@tmpb d%1,5-quinone (D)
175 \decalinev[begj]{#2}%
176 \else\if\@tmpb e%1,7-quinone (E)
177 \decalinev[bdfi]{#2}%
178 \else\if\@tmpb f%1,7-quinone (F)
179 \decalinev[cegj]{#2}%
180 \else\if\@tmpb g%1,7-quinone (G)
181 \decalinev[behj]{#2}%
182 \else\if\@tmpb h%1,7-quinone (H)
183 \decalinev[adgi]{#2}%
184 \fi\fi\fi\fi\fi\fi\fi\fi\fi%
185 \else\if\@tmpa P%quinone having 4 carbonyl groups
186 \if\@tmpb a% (A)1,4,5,8-type
187 \decalinev[bgk]{#2}%
188 \else\ifx\@tmpb\empty% (A)1,4,5,8-type
189 \decalinev[bgk]{#2}%
190 \else\if\@tmpb b% (B)1,2,5,8-type
191 \decalinev[cgk]{#2}%
192 \fi\fi\fi%
193 \else\if\@tmpa Q%1,2,3,4
194 \decalinev[fhk]{#2}%
195 \else\if\@tmpa O%quinone having 4 carbonyl groups
196 \if\@tmpb a% (A)1,2,5,6-type
197 \decalinev[chk]{#2}%
198 \else\ifx\@tmpb\empty% (A)1,2,5,6-type
199 \decalinev[chk]{#2}%
200 \else\if\@tmpb b% (B)1,2,7,8-type
201 \decalinev[cfk]{#2}%
202 \else\if\@tmpb c% (C)1,2,3,5-type
203 \decalinev[dgi]{#2}%
204 \else\if\@tmpb d% (D)1,2,3,7-type
205 \decalinev[dfi]{#2}%
206 \fi\fi\fi\fi\fi%
207 \else% added 1998/11/23 by S. Fujita
208 \decalinev[@#1]{#2}%
209 \fi\fi\fi\fi\fi\fi\fi\fi%
210 }%               %end of \naphdrv macro

7.2 Vertical-bottom type

The macro \naphdrvb is used for drawing naphthalene derivatives of verticalbottom type as well as naphthoquinone derivatives of vertical-bottom type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \decalinevb.

The macro \naphdrvb has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

 **********************************************
 * naphthalene and naphthoquinone derivatives *
 *  (vertical-bottom type)                    *
 **********************************************
   \naphdrvb[OPT]{SUBSLIST}

The OPT argument contains one or two characters.

The character ‘A’ is used for designating an aromatic bond pattern, in which each of the six-membered rings has a central circle.

     OPT (bond pattern)
           none       :  naphthalene
           A          :  aromatic circle

One character ‘p’ or two characters (p?) in the OPT argument designate a pquinone nucleus, which is present in the right-hand ring for the case of ‘p’ or ‘pa’ and in the left-hand ring for the case of ‘pb’.

           p or pa    :  1,4-quinone (A) left aromatic, right quinone
           pb         :  1,4-quinone (B) right aromatic, left quinone

One character ‘o’ or two characters (o?) in the OPT argument designate an o-quinone nucleus in either side of the fused ring. The second character of each argument indicates one of the possible type of bond patterns.

           o or oa    :  o-quinone (A)
           ob         :  o-quinone (B)
           oc         :  o-quinone (C)
           od         :  o-quinone (D)
           oe         :  o-quinone (E)
           of         :  o-quinone (F)

One character ‘q’ or two characters (q?) in the OPT argument designate an 2,6-, 1,5- or 1,7-quinone nucleus. The second character of each argument indicates one of the possible type of bond patterns.

           q or qa    :  2,6-quinone (A)
           qb         :  2,6-quinone (B)  3,7
           qc         :  1,5-quinone (C)
           qd         :  1,5-quinone (D)  4,8
           qe         :  1,7-quinone (E)
           qf         :  1,7-quinone (F)  2,8
           qg         :  1,7-quinone (G)  4,6
           qh         :  1,7-quinone (H)  3,5

One character ‘P’ or two characters (P?) in the OPT argument designate a 1,4,5,8- or 1,2,5,8-quinone nucleus.

           P or Pa    :  1,4,5,8-quinone (A)
           Pb         :  1,2,5,8-quinone (B)

One character ‘Q’ or ‘O’ or two characters (O?) in the OPT argument designate quinone nuclei of remaining types.

           Q          :  1,2,3,4-quinone
           O or Oa    :  1,2,5,6-quinone (A)
           Ob         :  1,2,7,8-quinone (B)
           Oc         :  1,2,3,5-quinone (C)
           Od         :  1,2,3,7-quinone (D)

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

     SUBSLIST: list of substituents
       for n = 1 to 8
           nD         :  exocyclic double bond at n-atom
           n or nS    :  exocyclic single bond at n-atom
           nA         :  alpha single bond at n-atom
           nB         :  beta single bond at n-atom
           nSA        :  alpha single bond at n-atom (boldface)
           nSB        :  beta single bond at n-atom (dotted line)

            nSa        :  alpha (not specified) single bond at n-atom
            nSb        :  beta (not specified) single bond at n-atom
            nSd        :  alpha single bond at n-atom (dotted line)
                            with an alternative direction to nSA
            nSu        :  beta single bond at n-atom (boldface)
                            with an alternative direction to nSB
            nFA        :  alpha single bond at n-atom (dotted line)
                            for ring fusion
            nFB        :  beta single bond at n-atom (boldface)
                            for ring fusion
            nGA        :  alpha single bond at n-atom (dotted line)
                            for the other ring fusion
            nGB        :  beta single bond at n-atom (boldface)
                            for the other ring fusion

        e.g.
         \naphdrvb{1==Cl;2==F}
         \naphdrvb[A]{1==Cl;4==F;2==CH$_{3}$}

\naphdrvb
211 \def\naphdrvb{\@ifnextchar[{\@naphdrvb}{\@naphdrvb[r]}}
212 \def\@naphdrvb[#1]#2{%
213 \iforigpt \typeout{command `naphdrvb' is based on `decalinevb'.}\fi%
214 \expandafter\threech@r#1{}{}%
215 \if\@tmpa r% right-handed set of double bonds
216 \decalinevb[acfhk]{#2}%
217 \else\if\@tmpa A%aromatic circle
218 \decalinevb[AB]{#2}%

219 %

220 \else\if\@tmpa p%p-quinone
221 \if\@tmpb a% (A)
222 \decalinevb[bfhk]{#2}%
223 \else\ifx\@tmpb\empty% (A)
224 \decalinevb[bfhk]{#2}%
225 \else\if\@tmpb b% (B)
226 \decalinevb[acgk]{#2}%
227 \fi\fi\fi%
228 \else%
229 \if\@tmpa o%o-quinone
230 \if\@tmpb a% (A)
231 \decalinevb[cfhk]{#2}%
232 \else\ifx\@tmpb\empty% (A)
233 \decalinevb[cfhk]{#2}%
234 \else\if\@tmpb b%o-quinone (B)
235 \decalinevb[dfhj]{#2}%
236 \else\if\@tmpb c%o-quinone (C)
237 \decalinevb[afhk]{#2}%

238 \else\if\@tmpb d%o-quinone (D)
239 \decalinevb[achk]{#2}%
240 \else\if\@tmpb e%o-quinone (E)
241 \decalinevb[acei]{#2}%
242 \else\if\@tmpb f%o-quinone (F)
243 \decalinevb[acfk]{#2}%
244 \fi\fi\fi\fi\fi\fi%
245 \fi%
246 %%%
247 \else\if\@tmpa q%various quinone
248 \if\@tmpb a% (A)2,6-type
249 \decalinevb[cehj]{#2}%
250 \else\ifx\@tmpb\empty% (A)2,6-type
251 \decalinevb[cehj]{#2}%
252 \else\if\@tmpb b%2,6-quinone (B)
253 \decalinevb[adfi]{#2}%
254 \else\if\@tmpb c%1,5-quinone (C)
255 \decalinevb[bdgi]{#2}%
256 \else\if\@tmpb d%1,5-quinone (D)
257 \decalinevb[begj]{#2}%
258 \else\if\@tmpb e%1,7-quinone (E)
259 \decalinevb[bdfi]{#2}%
260 \else\if\@tmpb f%1,7-quinone (F)
261 \decalinevb[cegj]{#2}%
262 \else\if\@tmpb g%1,7-quinone (G)
263 \decalinevb[behj]{#2}%
264 \else\if\@tmpb h%1,7-quinone (H)
265 \decalinevb[adgi]{#2}%
266 \fi\fi\fi\fi\fi\fi\fi\fi\fi%
267 \else\if\@tmpa P%quinone having 4 carbonyl groups
268 \if\@tmpb a% (A)1,4,5,8-type
269 \decalinevb[bgk]{#2}%
270 \else\ifx\@tmpb\empty% (A)1,4,5,8-type
271 \decalinevb[bgk]{#2}%
272 \else\if\@tmpb b% (B)1,2,5,8-type
273 \decalinevb[cgk]{#2}%
274 \fi\fi\fi%
275 \else\if\@tmpa Q%1,2,3,4
276 \decalinevb[fhk]{#2}%
277 \else\if\@tmpa O%quinone having 4 carbonyl groups
278 \if\@tmpb a% (A)1,2,5,6-type
279 \decalinevb[chk]{#2}%
280 \else\ifx\@tmpb\empty% (A)1,2,5,6-type
281 \decalinevb[chk]{#2}%
282 \else\if\@tmpb b% (B)1,2,7,8-type
283 \decalinevb[cfk]{#2}%
284 \else\if\@tmpb c% (C)1,2,3,5-type
285 \decalinevb[dgi]{#2}%
286 \else\if\@tmpb d% (D)1,2,3,7-type
287 \decalinevb[dfi]{#2}%

288 \fi\fi\fi\fi\fi%
289 \else% added 1998/11/23 by S. Fujita
290 \decalinevb[@#1]{#2}%
291 \fi\fi\fi\fi\fi\fi\fi\fi%
292 }%               %end of \naphdrvb macro

7.3 Vertical-top type

The macro \naphdrvt is used for drawing naphthalene derivatives of vertical-top type as well as naphthoquinone derivatives of vertical-top type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \decalinevt.

The macro \naphdrvt has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

 **********************************************
 * naphthalene and naphthoquinone derivatives *
 *  (vertical-top type)                       *
 **********************************************
   \naphdrvt[OPT]{SUBSLIST}

The OPT argument contains one or two characters.

The character ‘A’ is used for designating an aromatic bond pattern, in which each of the six-membered rings has a central circle.

     OPT (bond pattern)
           none       :  naphthalene
           A          :  aromatic circle

One character ‘p’ or two characters (p?) in the OPT argument designate a pquinone nucleus, which is present in the right-hand ring for the case of ‘p’ or ‘pa’ and in the left-hand ring for the case of ‘pb’.

           p or pa    :  1,4-quinone (A) left aromatic, right quinone
           pb         :  1,4-quinone (B) right aromatic, left quinone

One character ‘o’ or two characters (o?) in the OPT argument designate an o-quinone nucleus in either side of the fused ring. The second character of each argument indicates one of the possible type of bond patterns.

           o or oa    :  o-quinone (A)
           ob         :  o-quinone (B)
           oc         :  o-quinone (C)

           od         :  o-quinone (D)
           oe         :  o-quinone (E)
           of         :  o-quinone (F)

One character ‘q’ or two characters (q?) in the OPT argument designate an 2,6-, 1,5- or 1,7-quinone nucleus. The second character of each argument indicates one of the possible type of bond patterns.

           q or qa    :  2,6-quinone (A)
           qb         :  2,6-quinone (B)  3,7
           qc         :  1,5-quinone (C)
           qd         :  1,5-quinone (D)  4,8
           qe         :  1,7-quinone (E)
           qf         :  1,7-quinone (F)  2,8
           qg         :  1,7-quinone (G)  4,6
           qh         :  1,7-quinone (H)  3,5

One character ‘P’ or two characters (P?) in the OPT argument designate a 1,4,5,8- or 1,2,5,8-quinone nucleus.

           P or Pa    :  1,4,5,8-quinone (A)
           Pb         :  1,2,5,8-quinone (B)

One character ‘Q’ or ‘O’ or two characters (O?) in the OPT argument designate quinone nuclei of remaining types.

           Q          :  1,2,3,4-quinone
           O or Oa    :  1,2,5,6-quinone (A)
           Ob         :  1,2,7,8-quinone (B)
           Oc         :  1,2,3,5-quinone (C)
           Od         :  1,2,3,7-quinone (D)

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

     SUBSLIST: list of substituents
       for n = 1 to 8
           nD         :  exocyclic double bond at n-atom

            n or nS    :  exocyclic single bond at n-atom
            nA         :  alpha single bond at n-atom
            nB         :  beta single bond at n-atom
            nSA        :  alpha single bond at n-atom (boldface)
            nSB        :  beta single bond at n-atom (dotted line)
            nSa        :  alpha (not specified) single bond at n-atom
            nSb        :  beta (not specified) single bond at n-atom
            nSd        :  alpha single bond at n-atom (dotted line)
                            with an alternative direction to nSA
            nSu        :  beta single bond at n-atom (boldface)
                            with an alternative direction to nSB
            nFA        :  alpha single bond at n-atom (dotted line)
                            for ring fusion
            nFB        :  beta single bond at n-atom (boldface)
                            for ring fusion
            nGA        :  alpha single bond at n-atom (dotted line)
                            for the other ring fusion
            nGB        :  beta single bond at n-atom (boldface)
                            for the other ring fusion

        e.g.
         \naphdrvt{1==Cl;2==F}
         \naphdrvt[A]{1==Cl;4==F;2==CH$_{3}$}

\naphdrvt
293 \def\naphdrvt{\@ifnextchar[{\@naphdrvt}{\@naphdrvt[r]}}
294 \def\@naphdrvt[#1]#2{%
295 \iforigpt \typeout{command `naphdrvt' is based on `decalinevt'.}\fi%
296 \expandafter\threech@r#1{}{}%
297 \if\@tmpa r% right-handed set of double bonds
298 \decalinevt[acfhk]{#2}%
299 \else\if\@tmpa A%aromatic circle
300 \decalinevt[AB]{#2}%

301 %

302 \else\if\@tmpa p%p-quinone
303 \if\@tmpb a% (A)
304 \decalinevt[bfhk]{#2}%
305 \else\ifx\@tmpb\empty% (A)
306 \decalinevt[bfhk]{#2}%
307 \else\if\@tmpb b% (B)
308 \decalinevt[acgk]{#2}%
309 \fi\fi\fi%
310 \else%
311 \if\@tmpa o%o-quinone
312 \if\@tmpb a% (A)
313 \decalinevt[cfhk]{#2}%
314 \else\ifx\@tmpb\empty% (A)

315 \decalinevt[cfhk]{#2}%
316 \else\if\@tmpb b%o-quinone (B)
317 \decalinevt[dfhj]{#2}%
318 \else\if\@tmpb c%o-quinone (C)
319 \decalinevt[afhk]{#2}%
320 \else\if\@tmpb d%o-quinone (D)
321 \decalinevt[achk]{#2}%
322 \else\if\@tmpb e%o-quinone (E)
323 \decalinevt[acei]{#2}%
324 \else\if\@tmpb f%o-quinone (F)
325 \decalinevt[acfk]{#2}%
326 \fi\fi\fi\fi\fi\fi%
327 \fi%
328 %%%
329 \else\if\@tmpa q%various quinone
330 \if\@tmpb a% (A)2,6-type
331 \decalinevt[cehj]{#2}%
332 \else\ifx\@tmpb\empty% (A)2,6-type
333 \decalinevt[cehj]{#2}%
334 \else\if\@tmpb b%2,6-quinone (B)
335 \decalinevt[adfi]{#2}%
336 \else\if\@tmpb c%1,5-quinone (C)
337 \decalinevt[bdgi]{#2}%
338 \else\if\@tmpb d%1,5-quinone (D)
339 \decalinevt[begj]{#2}%
340 \else\if\@tmpb e%1,7-quinone (E)
341 \decalinevt[bdfi]{#2}%
342 \else\if\@tmpb f%1,7-quinone (F)
343 \decalinevt[cegj]{#2}%
344 \else\if\@tmpb g%1,7-quinone (G)
345 \decalinevt[behj]{#2}%
346 \else\if\@tmpb h%1,7-quinone (H)
347 \decalinevt[adgi]{#2}%
348 \fi\fi\fi\fi\fi\fi\fi\fi\fi%
349 \else\if\@tmpa P%quinone having 4 carbonyl groups
350 \if\@tmpb a% (A)1,4,5,8-type
351 \decalinevt[bgk]{#2}%
352 \else\ifx\@tmpb\empty% (A)1,4,5,8-type
353 \decalinevt[bgk]{#2}%
354 \else\if\@tmpb b% (B)1,2,5,8-type
355 \decalinevt[cgk]{#2}%
356 \fi\fi\fi%
357 \else\if\@tmpa Q%1,2,3,4
358 \decalinevt[fhk]{#2}%
359 \else\if\@tmpa O%quinone having 4 carbonyl groups
360 \if\@tmpb a% (A)1,2,5,6-type
361 \decalinevt[chk]{#2}%
362 \else\ifx\@tmpb\empty% (A)1,2,5,6-type
363 \decalinevt[chk]{#2}%
364 \else\if\@tmpb b% (B)1,2,7,8-type

365 \decalinevt[cfk]{#2}%
366 \else\if\@tmpb c% (C)1,2,3,5-type
367 \decalinevt[dgi]{#2}%
368 \else\if\@tmpb d% (D)1,2,3,7-type
369 \decalinevt[dfi]{#2}%
370 \fi\fi\fi\fi\fi%
371 \else% added 1998/11/23 by S. Fujita
372 \decalinevt[@#1]{#2}%
373 \fi\fi\fi\fi\fi\fi\fi\fi%
374 }%               %end of \naphdrvt macro

7.4 Horizontal type

The macro \naphdrh is used for drawing naphthalene derivatives of horizontal type as well as naphtoquinone derivatives of horizontal type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \decalineh.

The macro \naphdrh has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

  **********************************************
  * naphthalene and naphthoquinone derivatives *
  *  (horizontal type)                         *
  **********************************************
    \naphdrh[OPT]{SUBSLIST}

The OPT argument contains one or two characters for designate aromatic bond patterns or for naphthoquinone bond patterns. The meaning of these characters is the same as described in the definition of the macro \naphdrv.

      OPT (bond pattern)
            none       :  naphthalene
            A          :  aromatic circle
            p or pa    :  1,4-quinone (A) left aromatic, right quinone
            pb         :  1,4-quinone (B) right aromatic, left quinone
            o or oa    :  o-quinone (A)
            ob         :  o-quinone (B)
            oc         :  o-quinone (C)
            od         :  o-quinone (D)
            oe         :  o-quinone (E)
            of         :  o-quinone (F)

q or qa    :  2,6-quinone (A)
qb         :  2,6-quinone (B)  3,7
qc         :  1,5-quinone (C)
qd         :  1,5-quinone (D)  4,8
qe         :  1,7-quinone (E)
qf         :  1,7-quinone (F)  2,8
qg         :  1,7-quinone (G)  4,6
qh         :  1,7-quinone (H)  3,5
P or Pa    :  1,4,5,8-quinone (A)
Pb         :  1,2,5,8-quinone (B)
Q          :  1,2,3,4-quinone
O or Oa    :  1,2,5,6-quinone (A)
Ob         :  1,2,7,8-quinone (B)
Oc         :  1,2,3,5-quinone (C)
Od         :  1,2,3,7-quinone (D)

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 8
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)
                for the other ring fusion
nGB        :  beta single bond at n-atom (boldface)

                            for the other ring fusion

        e.g.
         \naphdrh{1==Cl;2==F}
         \naphdrh[A]{1==Cl;4==F;2==CH$_{3}$}

\naphdrh
375 \def\naphdrh{\@ifnextchar[{\@naphdrh}{\@naphdrh[r]}}
376 \def\@naphdrh[#1]#2{%
377 \iforigpt \typeout{command `nathdrh' is based on `decalineh'.}\fi%
378 \expandafter\threech@r#1{}{}%
379 \if\@tmpa r% right-handed set of double bonds
380 \decalineh[acfhk]{#2}%
381 \else\if\@tmpa A%aromatic circle
382 \decalineh[AB]{#2}%

383 %

384 \else\if\@tmpa p%p-quinone
385 \if\@tmpb a% (A)
386 \decalineh[bfhk]{#2}%
387 \else\ifx\@tmpb\empty% (A)
388 \decalineh[bfhk]{#2}%
389 \else\if\@tmpb b% (B)
390 \decalineh[acgk]{#2}%
391 \fi\fi\fi%
392 \else%
393 \if\@tmpa o%o-quinone
394 \if\@tmpb a% (A)
395 \decalineh[cfhk]{#2}%
396 \else\ifx\@tmpb\empty% (A)
397 \decalineh[cfhk]{#2}%
398 \else\if\@tmpb b%o-quinone (B)
399 \decalineh[dfhj]{#2}%
400 \else\if\@tmpb c%o-quinone (C)
401 \decalineh[afhk]{#2}%
402 \else\if\@tmpb d%o-quinone (D)
403 \decalineh[achk]{#2}%
404 \else\if\@tmpb e%o-quinone (E)
405 \decalineh[acei]{#2}%
406 \else\if\@tmpb f%o-quinone (F)
407 \decalineh[acfk]{#2}%
408 \fi\fi\fi\fi\fi\fi%
409 \fi%
410 % %%
411 \else\if\@tmpa q%various quinone
412 \if\@tmpb a% (A)2,6-type
413 \decalineh[cehj]{#2}%
414 \else\ifx\@tmpb\empty% (A)2,6-type

415 \decalineh[cehj]{#2}%
416 \else\if\@tmpb b%2,6-quinone (B)
417 \decalineh[adfi]{#2}%
418 \else\if\@tmpb c%1,5-quinone (C)
419 \decalineh[bdgi]{#2}%
420 \else\if\@tmpb d%1,5-quinone (D)
421 \decalineh[begj]{#2}%
422 \else\if\@tmpb e%1,7-quinone (E)
423 \decalineh[bdfi]{#2}%
424 \else\if\@tmpb f%1,7-quinone (F)
425 \decalineh[cegj]{#2}%
426 \else\if\@tmpb g%1,7-quinone (G)
427 \decalineh[behj]{#2}%
428 \else\if\@tmpb h%1,7-quinone (H)
429 \decalineh[adgi]{#2}%
430 \fi\fi\fi\fi\fi\fi\fi\fi\fi%
431 \else\if\@tmpa P%quinone having 4 carbonyl groups
432 \if\@tmpb a% (A)1,4,5,8-type
433 \decalineh[bgk]{#2}%
434 \else\ifx\@tmpb\empty% (A)1,4,5,8-type
435 \decalineh[bgk]{#2}%
436 \else\if\@tmpb b% (B)1,2,5,8-type
437 \decalineh[cgk]{#2}%
438 \fi\fi\fi%
439 \else\if\@tmpa Q%1,2,3,4
440 \decalineh[fhk]{#2}%
441 \else\if\@tmpa O%quinone having 4 carbonyl groups
442 \if\@tmpb a% (A)1,2,5,6-type
443 \decalineh[chk]{#2}%
444 \else\ifx\@tmpb\empty% (A)1,2,5,6-type
445 \decalineh[chk]{#2}%
446 \else\if\@tmpb b% (B)1,2,7,8-type
447 \decalineh[cfk]{#2}%
448 \else\if\@tmpb c% (C)1,2,3,5-type
449 \decalineh[dgi]{#2}%
450 \else\if\@tmpb d% (D)1,2,3,7-type
451 \decalineh[dfi]{#2}%
452 \fi\fi\fi\fi\fi%
453 \else% added 1998/11/23 by S. Fujita
454 \decalineh[@#1]{#2}%
455 \fi\fi\fi\fi\fi\fi\fi\fi%
456 }%               %end of \naphdrh macro

8 Tetraline derivatives

8.1 Vertical type

The macro \tetralinev is used for drawing naphthalene derivatives of vertical type as well as naphthoquinone derivatives of vertical type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \decalinev.

The macro \tetralinev has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

*************************
* tetraline derivatives *
*  (vertical type)      *
*************************
\tetralinev[OPT]{SUBSLIST}

The OPT argument contains one or two characters for designate aromatic bond patterns or for naphthoquinone bond patterns.

OPT = none       :  tetraline
A          :  aromatic circle
e or ea    :  1,2-double bond
eb         :  2,3-double bond
ec         :  3,4-double bond

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 4
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)

                            for ring fusion
            nFB        :  beta single bond at n-atom (boldface)
                            for ring fusion
            nGA        :  alpha single bond at n-atom (dotted line)
                            for the other ring fusion
            nGB        :  beta single bond at n-atom (boldface)
                            for the other ring fusion
         for n=5 or 8
            nSa        :  bias bond
                         This option can be used to avoid over-crowding
                         betweem 1 and 8 or between 4 and 5.

        e.g.
         \tetralinev{1==Cl;2==F}
         \tetralinev[A]{1==Cl;4==F;2==CH$_{3}$}
         \tetralinev[eb]{1D==O;4SA==MeO;4SB==OMe;5W==Cl;6==Cl;7==Cl;8==Cl}

\tetralinev
457 \def\tetralinev{\@ifnextchar[{\@tetralinev}{\@tetralinev[r]}}
458 \def\@tetralinev[#1]#2{%
459 \iforigpt \typeout{command `tetralinev' is based on `decalinev'.}\fi%
460 \expandafter\threech@r#1{}{}%
461 \if\@tmpa r% right-handed set of double bonds
462  \decalinev[fhk]{#2}%
463 \else\if\@tmpa A%aromatic circle
464  \decalinev[A]{#2}%
465 \else\if\@tmpa e%inner double bond
466 \if\@tmpb a% (A)%
467  \decalinev[afhk]{#2}%
468 \else\ifx\@tmpb\empty% (A)
469  \decalinev[afhk]{#2}%
470 \else\if\@tmpb b% (B)
471  \decalinev[bfhk]{#2}%
472 \else\if\@tmpb c% (C)
473  \decalinev[cfhk]{#2}%
474 \fi\fi\fi\fi%
475 \else%1998/11/23 by Shinsaku Fujita
476  \decalinev[@#1]{#2}%
477 \fi\fi\fi%
478 }%               %end of \tetralinev macro

8.2 Vertical-bottom type

The macro \tetralinevb is used for drawing naphthalene derivatives of verticalbottom type as well as naphthoquinone derivatives of vertical type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \decalinevb.

The macro \tetralinevb has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

********************************
* tetraline derivatives        *
*  (vertical-bottom type)      *
********************************
\tetralinevb[OPT]{SUBSLIST}

The OPT argument contains one or two characters for designate aromatic bond patterns or for naphthoquinone bond patterns.

OPT = none       :  tetraline
A          :  aromatic circle
e or ea    :  1,2-double bond
eb         :  2,3-double bond
ec         :  3,4-double bond

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 4
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)

                            for ring fusion
            nFB        :  beta single bond at n-atom (boldface)
                            for ring fusion
            nGA        :  alpha single bond at n-atom (dotted line)
                            for the other ring fusion
            nGB        :  beta single bond at n-atom (boldface)
                            for the other ring fusion
         for n=5 or 8
            nSa        :  bias bond
                         This option can be used to avoid over-crowding
                         betweem 1 and 8 or between 4 and 5.

        e.g.
         \tetralinevb{1==Cl;2==F}
         \tetralinevb[A]{1==Cl;4==F;2==CH$_{3}$}
         \tetralinevb[eb]{1D==O;4SA==MeO;4SB==OMe;5W==Cl;6==Cl;7==Cl;8==Cl}

\tetralinevb
479 \def\tetralinevb{\@ifnextchar[{\@tetralinevb}{\@tetralinevb[r]}}
480 \def\@tetralinevb[#1]#2{%
481 \iforigpt \typeout{command `tetralinevb' is based on `decalinevb'.}\fi%
482 \expandafter\threech@r#1{}{}%
483 \if\@tmpa r% right-handed set of double bonds
484  \decalinevb[fhK]{#2}%  %[fhK] in place of [fhk]
485 \else\if\@tmpa A%aromatic circle
486  \decalinevb[A]{#2}%
487 \else\if\@tmpa e%inner double bond
488 \if\@tmpb a% (A)%
489  \decalinevb[afhK]{#2}%
490 \else\ifx\@tmpb\empty% (A)
491  \decalinevb[afhK]{#2}%
492 \else\if\@tmpb b% (B)
493  \decalinevb[bfhK]{#2}%
494 \else\if\@tmpb c% (C)
495  \decalinevb[cfhK]{#2}%
496 \fi\fi\fi\fi%
497 \else%1998/11/23 by Shinsaku Fujita
498  \decalinevb[@#1]{#2}%
499 \fi\fi\fi%
500 }%               %end of \tetralinevb macro

8.3 Vertical-top type

The macro \tetralinevt is used for drawing naphthalene derivatives of verticaltop type as well as naphthoquinone derivatives of vertical type. The skeleton and endocyclic double bonds are drawn directly, while substituents and exocyclic bonds are placed by useing the inner macro \decalinevt.

The macro \tetralinevt has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

********************************
* tetraline derivatives        *
*  (vertical-bottom type)      *
********************************
\tetralinevt[OPT]{SUBSLIST}

The OPT argument contains one or two characters for designate aromatic bond patterns or for naphthoquinone bond patterns.

OPT = none       :  tetraline
A          :  aromatic circle
e or ea    :  1,2-double bond
eb         :  2,3-double bond
ec         :  3,4-double bond

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 4
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)

                            for ring fusion
            nFB        :  beta single bond at n-atom (boldface)
                            for ring fusion
            nGA        :  alpha single bond at n-atom (dotted line)
                            for the other ring fusion
            nGB        :  beta single bond at n-atom (boldface)
                            for the other ring fusion
         for n=5 or 8
            nSa        :  bias bond
                         This option can be used to avoid over-crowding
                         betweem 1 and 8 or between 4 and 5.

        e.g.
         \tetralinevt{1==Cl;2==F}
         \tetralinevt[A]{1==Cl;4==F;2==CH$_{3}$}
         \tetralinevt[eb]{1D==O;4SA==MeO;4SB==OMe;5W==Cl;6==Cl;7==Cl;8==Cl}

\tetralinevt
501 \def\tetralinevt{\@ifnextchar[{\@tetralinevt}{\@tetralinevt[r]}}
502 \def\@tetralinevt[#1]#2{%
503 \iforigpt \typeout{command `tetralinevt' is based on `decalinevt'.}\fi%
504 \expandafter\threech@r#1{}{}%
505 \if\@tmpa r% right-handed set of double bonds
506  \decalinevt[fhk]{#2}%
507 \else\if\@tmpa A%aromatic circle
508  \decalinevt[A]{#2}%
509 \else\if\@tmpa e%inner double bond
510 \if\@tmpb a% (A)%
511  \decalinevt[afhk]{#2}%
512 \else\ifx\@tmpb\empty% (A)
513  \decalinevt[afhk]{#2}%
514 \else\if\@tmpb b% (B)
515  \decalinevt[bfhk]{#2}%
516 \else\if\@tmpb c% (C)
517  \decalinevt[cfhk]{#2}%
518 \fi\fi\fi\fi%
519 \else%1998/11/23 by Shinsaku Fujita
520  \decalinevt[@#1]{#2}%
521 \fi\fi\fi%
522 }%               %end of \tetralinevt macro

8.4 Horizontal type

The macro \tetralineh is used for drawing tetraline derivatives of horizontal type.

The macro \tetralineh has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

*************************
* tetraline derivatives *
*  (horizontal type)    *
*************************
\tetralineh[OPT]{SUBSLIST}

The OPT argument contains one or two characters for designate aromatic bond patterns or for naphthoquinone bond patterns.

OPT = none       :  tetraline
A          :  aromatic circle
e or ea    :  1,2-double bond
eb         :  2,3-double bond
ec         :  3,4-double bond

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 4
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)

                            for the other ring fusion
            nGB        :  beta single bond at n-atom (boldface)
                            for the other ring fusion
         for n=5 or 8
            nSa        :  bias bond
                         This option can be used to avoid over-crowding
                         betweem 1 and 8 or between 4 and 5.

        e.g.
         \tetralineh{1==Cl;2==F}
         \tetralineh[A]{1==Cl;4==F;2==CH$_{3}$}
         \tetralineh[eb]{1D==O;4SA==MeO;4SB==OMe;5W==Cl;6==Cl;7==Cl;8==Cl}

\tetralineh
523 \def\tetralineh{\@ifnextchar[{\@tetralineh}{\@tetralineh[r]}}
524 \def\@tetralineh[#1]#2{%
525 \iforigpt \typeout{command `tetralineh' is based on `decalineh'.}\fi%
526 \expandafter\threech@r#1{}{}%
527 \if\@tmpa r% right-handed set of double bonds
528  \decalineh[fhk]{#2}%
529 \else\if\@tmpa A%aromatic circle
530  \decalineh[A]{#2}%
531 \else\if\@tmpa e%inner double bond
532 \if\@tmpb a% (A)%
533  \decalineh[afhk]{#2}%
534 \else\ifx\@tmpb\empty% (A)
535  \decalineh[afhk]{#2}%
536 \else\if\@tmpb b% (B)
537  \decalineh[bfhk]{#2}%
538 \else\if\@tmpb c% (C)
539  \decalineh[cfhk]{#2}%
540 \fi\fi\fi\fi%
541 \else% added 1998/11/23 by S. Fujita
542 \decalineh[@#1]{#2}%
543 \fi\fi\fi%
544 }%               %end of \tetralineh macro

9 Anthracene derivatives

9.1 Perhydro-anthracene derivatives of Vertical type

The macro \hanthracenev is used for drawing perhydroanthracene derivatives of vertical type.

The macro \hanthracenev has an argument SUBSLIST as well as an optional argument BONDLIST.

***********************************
* perhydro anthracene derivatives *
*  (vertical type)                *
***********************************
\hanthracenev[BONDLIST]{SUBSLIST}

The BONDLIST argument contains one or more characters selected form a to p, each of which indicates the presence of an inner (endcyclic) double bond on the corresponding position. The character A, B, or C indicates an aromatic bond pattern (circle).

BONDLIST =
none       :  hanthracene
a          :  1,2-double bond
b          :  2,3-double bond
c          :  3,4-double bond
d          :  4,4a-double bond
e          :  10,4a-double bond
f          :  10,10a-double bond
g          :  5,10a-double bond
h          :  5,6-double bond
i          :  6,7-double bond
j          :  8,7-double bond
k          :  8,8a-double bond
l          :  9,8a-double bond
m          :  9,9a-double bond
n          :  1,9a-double bond
o          :  4a,9a-double bond
O          :  4a,9a-double bond in the other ring
p          :  10a,8a-double bond
P          :  10a,8a-double bond in the other ring
A          :  right aromatic circle
B          :  central aromatic circle
C          :  left aromatic circle

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents

nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)
                for the other ring fusion
nGB        :  beta single bond at n-atom (boldface)
                for the other ring fusion

for 11 (fused positions, 9a and 4a)
11FA       :  alpha single bond at 9a
11FB       :  beta single bond at 9a
11FU       :  unspecified single bond at 9a
11GA       :  alpha single bond at 4a
11GB       :  beta single bond at 4a
11GU       :  unspecified single bond at 4a

for 12 (fused positions, 8a and 10a)
12FA       :  alpha single bond at 8a
12FB       :  beta single bond at 8a
12FU       :  unspecified single bond at 8a
12GA       :  alpha single bond at 10a
12GB       :  beta single bond at 10a
12GU       :  unspecified single bond at 10a

e.g.
\hanthracenev{1==Cl;2==F;{{10}}==OH}

         \hanthracenev[c]{1==Cl;4==F;2==CH$_{3}$;
             {{11}FA}==OH;{{11}FB}==OH}
         \hanthracenev[eb]{1D==O;4SA==MeO;4SB==OMe;5W==Cl;6==Cl;7==Cl;8==Cl}

\hanthracenev
545 \def\hanthracenev{\@ifnextchar[{\@hanthracenev[@}{\@hanthracenev[r]}}
546 \def\@hanthracenev[#1]#2{%
547 % bonds for outer skeleton
548 \@reset@ylsw\reset@@yl
549 \ylposition{#2}{-7}{0}{2}{0}%  %for 8 positions
550 \if@ylsw \ifx\@@ylii\empty
551 \def\@@ylii{0}\def\@@yli{0}\fi
552 \else\ylposition{#2}{-1}{3}{7}{0}% for 5 to 7 positions
553 \fi
554 \if@ylsw\ifx\@@ylii\empty
555 \def\@@ylii{0}\def\@@yli{0}\fi\else
556 \ylposition{#2}{0}{0}{5}{0}% for 1 to 4 positions
557 \fi
558 \if@ylsw\ifx\@@ylii\empty
559 \def\@@ylii{-684}\def\@@yli{0}\fi\else
560 \ylposition{#2}{-8}{0}{2}{0}%for 9-position
561 \fi
562 \if@ylsw\ifx\@@ylii\empty
563 \def\@@ylii{-342}\def\@@yli{0}\fi\else
564 \ylposition{#2}{-6}{3}{5}{0}%for 10-position
565 \fi
566 \if@ylsw\ifx\@@ylii\empty
567 \def\@@ylii{-342}\def\@@yli{0}\fi\else
568 \ylfusedposition{#2}{-11}{-1}{1}{0}% for 9a and 4a
569 \fi
570 \if@ylsw\ifx\@@ylii\empty
571 \def\@@ylii{-342}\def\@@yli{0}\fi\else
572 \ylfusedposition{#2}{-12}{-1}{1}{0}% for 8a and 10a
573 \fi
574 \if@ylsw\ifx\@@ylii\empty
575 \def\@@ylii{0}\def\@@yli{0}\fi\fi
576 \if@ylsw
577 \yl@shiftii=\@ylii
578 \ifx\@@ylii\empty\else
579 \advance\yl@shiftii\@@ylii\fi
580 \yl@shifti=\@yli
581 \ifx\@@yli\empty\else
582 \advance\yl@shifti\@@yli\fi
583 \advance\yl@shiftii\yl@xdiff
584 \advance\yl@shifti\yl@ydiff
585 \begin{picture}(0,0)(-\yl@shiftii,-\yl@shifti)
586 \reset@yl@xydiff%1999/1/6 by S. Fujita
587 \else
588 \begin{picture}(1200,880)(-\shiftii,-\shifti)

589 \iforigpt \put(-\shiftii,-\shifti){\circle*{50}}%
590  \put(-\noshift,-\noshift){\circle{50}}%
591 \typeout{command `hanthracenev' origin: %
592  (\the\noshift,\the\noshift) ---> (\the\shiftii,\the\shifti)}\fi%
593 \fi
594 \put(0,406){\line(-5,-3){171}}%       %bond 8-7
595 \put(0,406){\line(5,-3){171}}%        %bond 8-8a
596 \put(0,0){\line(-5,3){171}}%          %bond 5-6
597 \put(0,0){\line(5,3){171}}%           %bond 5-10a
598 \put(171,103){\line(0,1){200}}%       %bond 10a-8a
599 \put(-171,103){\line(0,1){200}}%      %bond 6-7
600 \put(342,406){\line(-5,-3){171}}%     %bond 9-8a
601 \put(342,406){\line(5,-3){171}}%      %bond 9-9a
602 \put(342,0){\line(-5,3){171}}%        %bond 10-10a
603 \put(342,0){\line(5,3){171}}%         %bond 10-4a
604 \put(513,103){\line(0,1){200}}%       %bond 4a-9a
605 %%%%
606 \put(684,406){\line(-5,-3){171}}%     %bond 1-9a    1-6
607 \put(684,406){\line(5,-3){171}}%      %bond 1-2     1-2
608 \put(855,103){\line(0,1){200}}%       %bond 3-2     3-2
609 \put(684,0){\line(5,3){171}}%         %bond 4-3     4-3
610 \put(684,0){\line(-5,3){171}}%        %bond 4-4a    4-5
611 % inner double bonds
612 \@tfor\member:=#1\do{%
613 \if\member r%no endcyclic double bonds
614 \else\if\member a%
615 \put(690,364){\line(5,-3){126}}%      %double bond 1-2
616 \else\if\member b%
617 \put(822,129){\line(0,1){148}}%       %double bond 3-2
618 \else\if\member c%
619 \put(690,42){\line(5,3){126}}%        %double bond 4-3
620 \else\if\member d%
621 \put(690,42){\line(-5,3){126}}%       %double bond 4-4a
622 \else\if\member e%
623 \put(348,42){\line(5,3){126}}%        %double bond 10-4a
624 \else\if\member f%
625 \put(348,42){\line(-5,3){126}}%       %double bond 10-10a
626 \else\if\member g%
627 \put(6,42){\line(5,3){126}}%          %double bond 5-10a
628 \else\if\member h%
629 \put(-6,42){\line(-5,3){126}}%        %double bond 5-6
630 \else\if\member i%
631 \put(-138,129){\line(0,1){148}}%      %double bond 6-7
632 \else\if\member j%
633 \put(-6,364){\line(-5,-3){126}}%      %double bond 8-7
634 \else\if\member k%
635 \put(6,364){\line(5,-3){126}}%        %double bond 8-8a
636 \else\if\member l%
637 \put(348,364){\line(-5,-3){126}}%     %double bond 9-8a
638 \else\if\member m%

639 \put(348,364){\line(5,-3){126}}%      %double bond 9-9a
640 \else\if\member n%
641 \put(690,364){\line(-5,-3){126}}%     %double bond 1-9a
642 \else\if\member o%
643 \put(544,129){\line(0,1){148}}%       %double bond 4a-9a
644 \else\if\member O%
645 \put(482,129){\line(0,1){148}}%       %double bond 4a-9a at the other ring
646 \else\if\member p%
647 \put(138,129){\line(0,1){148}}%       %double bond 10a-8a
648 \else\if\member P%
649 \put(204,129){\line(0,1){148}}%       %double bond 10a-8a at the other ring
650 \else\if\member C%left aromatic circle
651 \put(0,203){\circle{240}}            %circle
652 \else\if\member B%right aromatic circle
653 \put(342,203){\circle{240}}%          %circle
654 \else\if\member A%right aromatic circle
655 \put(684,203){\circle{240}}%          %circle
656 \else
657 %setting fused rings
658 \expandafter\twoCH@R\member//%
659 \set@fusion@hanthrav
660 \fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi}
661 %exocyclic bonds and substituents
662 \put(0,0){\setsixringv{#2}{-7}{0}{2}{0}}%  %for 8 positions
663 \put(0,0){\setsixringv{#2}{-1}{3}{7}{0}}%  %for 5 to 7 positions
664 \put(684,0){\setsixringv{#2}{0}{0}{5}{0}}% %for 1 to 4 positions
665 \put(342,0){\setsixringv{#2}{-8}{0}{2}{0}}%for 9-position
666 \put(342,0){\setsixringv{#2}{-6}{3}{5}{0}}%for 10-position
667 %bonds and substituents for fused positions
668 \put(0,0){\setfusedbond{#2}{-12}{-1}{1}{0}}% for 8a and 10a
669 \put(342,0){\setfusedbond{#2}{-11}{-1}{1}{0}}% for 9a and 4a
670 \end{picture}}%               %end of \hanthracenev macro%

9.2 Anthracene and anthraquinone derivatives

The macro \anthracenev is used for drawing anthracene derivatives of vertical type.

The macro \anthracenev has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

  ********************************************
  * anthracene and anthraquinone derivatives *
  *  (vertical type)                         *
  ********************************************
    \anthracenev[OPT]{SUBSLIST}

The OPT argument contains one or two characters for designate aromatic bond patterns or for naphthoquinone bond patterns.

OPT (bond pattern)
none or r  :  right-handed double bonds
l          :  left-handed double bonds
A          :  aromatic circle
p or pa    :  9,10-anthraquinone (A)
pA         :  9,10-anthraquinone (circle type)
o          :  1,2-anthraquinone (A)
oa         :  1,2-anthraquinone (A')
oA         :  1,2-anthraquinone (circle type)
ob         :  2,3-antharquinone (B)
oc         :  1,2-anthraquinone (C)
q          :  1,4-anthraquinone (A)
qa         :  1,4-anthraquinone (A')
qA         :  1,4-anthraquinone (circle type)

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 10
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)

        e.g.
         \anthracenev{1==Cl;2==F;0FA==;0GA==}
         \anthracenev[A]{1==Cl;4==F;2==CH$_{3}$}
         \anthracenev[eb]{1D==O;4SA==MeO;4SB==OMe;%
              5W==Cl;6==Cl;7==Cl;8==Cl}

\anthracenev
671 \def\anthracenev{\@ifnextchar[{\@anthracenev}{\@anthracenev[r]}}
672 \def\@anthracenev[#1]#2{%
673 \iforigpt \typeout{command `anthracenev' is %
674  based on `hanthracenev'.}\fi%
675 %% \expandafter\threech@r#1{}{}%error1998/11/23 by Shinsaku Fujita
676 \expandafter\twoCH@R#1//%
677 \if\@tmpa r% right-handed set of double bonds
678  \hanthracenev[acflhjo]{#2}%
679 \else\if\@tmpa l% left-handed set of double bonds
680  \hanthracenev[acehjmp]{#2}%
681 \else\if\@tmpa A%aromatic circle
682  \hanthracenev[ABC]{#2}%
683 \else\if\@tmpa p%9,10-quinone
684 \if\@tmpb a% (A)%
685  \hanthracenev[achjop]{#2}%
686 \else\ifx\@tmpb\empty% (A)
687  \hanthracenev[achjop]{#2}%
688 \else\if\@tmpb A% circle type
689  \hanthracenev[AC]{#2}%
690 \else
691  \hanthracenev[achjop]{#2}%
692 \fi\fi\fi%
693 \else\if\@tmpa o%o-quinone
694 \if\@tmpb a% (A)%
695  \hanthracenev[cfjhlo]{#2}%
696 \else\ifx\@tmpb\empty% (A)
697  \hanthracenev[cehjmp]{#2}%
698 \else\if\@tmpb b% (B)
699  \hanthracenev[dfhjln]{#2}%
700 \else\if\@tmpb c% (C)
701  \hanthracenev[aehjmp]{#2}%
702 \else\if\@tmpb A% cicle type
703  \hanthracenev[cBC]{#2}%
704 \fi\fi\fi\fi\fi%
705 \else\if\@tmpa q%1,4-quinone
706 \if\@tmpb a% (A)%

707  \hanthracenev[bfhjlo]{#2}%
708 \else\ifx\@tmpb\empty% (A)
709  \hanthracenev[behjmp]{#2}%
710 \else\if\@tmpb A% circle type
711  \hanthracenev[AB]{#2}%
712 \else%
713  \hanthracenev[behjmp]{#2}%
714 \fi\fi\fi%
715 \else
716  \hanthracenev[@#1]{#2}%
717 \fi\fi\fi\fi\fi\fi%
718 }%               %end of \anthracenev macro%

10 Phenanthrene derivatives

10.1 Perhydro-phenanthrene derivatives of Vertical type

The macro \hphenanthrenev is used for drawing perhydrophenanthrene derivatives of vertical type.

The macro \hphenanthrenev has an argument SUBSLIST as well as an optional argument BONDLIST.

  *************************************
  * perhydro phenanthrene derivatives *
  *  (vertical type)                  *
  *************************************
    \hphenanthrenev[BONDLIST]{SUBSLIST}

The BONDLIST argument contains one or more characters selected form a to p, each of which indicates the presence of an inner (endcyclic) double bond on the corresponding position. The character A, B, or C indicates an aromatic bond pattern (circle).

      BONDLIST =
            none       :  hphenanthrene
            a          :  1,2-double bond
            b          :  2,3-double bond
            c          :  3,4-double bond
            d          :  4,4a-double bond
            e          :  4a,4b-double bond
            f          :  4b,5-double bond
            g          :  5,6-double bond
            h          :  6,7-double bond
            i          :  7,8-double bond
            j          :  8,8a-double bond

k          :  8a,9-double bond
l          :  9,10-double bond
m          :  10,10a-double bond
n          :  1,10a-double bond
o          :  4a,10a-double bond
O          :  4a,10a-double bond in the other ring
p          :  4b,8a-double bond
P          :  4b,8a-double bond in the other ring
A          :  right aromatic circle
B          :  central aromatic circle
C          :  left aromatic circle

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 10
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom
nB         :  beta single bond at n-atom
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion
nGA        :  alpha single bond at n-atom (dotted line)
                for the other ring fusion
nGB        :  beta single bond at n-atom (boldface)
                for the other ring fusion
for 11 (fused positions, 4a and 10a)
11FA       :  alpha single bond at 4a
11FB       :  beta single bond at 4a
11FC       :  alpha single bond (downward) at 4a
11FD       :  beta single bond (downward) at 4a

            11FU       :  unspecified single bond at 4a
            11GA       :  alpha single bond at 10a
            11GB       :  beta single bond at 10a
            11FC       :  alpha single bond (upward) at 10a
            11FD       :  beta single bond (upward) at 10a
            11GU       :  unspecified single bond at 10a
         for 12 (fused positions, 4b and 8a)
            12FA       :  alpha single bond at 4b
            12FB       :  beta single bond at 4b
            12FU       :  unspecified single bond at 4b
            12GA       :  alpha single bond at 8a
            12GB       :  beta single bond at 8a
            12GU       :  unspecified single bond at 8a

        e.g.
         \hphenanthrenev{1==Cl;2==F;{{10}}==OH}
         \hphenanthrenev[c]{1==Cl;4==F;2==CH$_{3}$;
             {{11}FA}==OH;{{11}FB}==OH}
         \hphenanthrenev[eb]{1D==O;4SA==MeO;4SB==OMe;
                    5W==Cl;6==Cl;7==Cl;8==Cl}

\hphenanthrenev
719 \def\hphenanthrenev{%
720 \@ifnextchar[{\@hphenanthrenev[@}{\@hphenanthrenev[r]}}
721 \def\@hphenanthrenev[#1]#2{%
722 % bonds for outer skeleton
723 \@reset@ylsw\reset@@yl
724 \ylposition{#2}{-4}{0}{2}{0}%   %for 5 position
725 \if@ylsw \ifx\@@ylii\empty
726 \def\@@ylii{0}\def\@@yli{0}\fi
727 \else
728 \ylposition{#2}{12}{3}{7}{-1}%   %for 6 to 8 positions
729 \fi
730 \if@ylsw\ifx\@@ylii\empty
731 \def\@@ylii{0}\def\@@yli{0}\fi\else
732 \ylposition{#2}{13}{2}{5}{-1}%   %for 9 to 10 positions
733 \fi
734 \if@ylsw\ifx\@@ylii\empty
735 \def\@@ylii{-342}\def\@@yli{0}\fi\else
736 \ylposition{#2}{4}{0}{4}{-1}% %for 1 to 3 positions
737 \fi
738 \if@ylsw\ifx\@@ylii\empty
739 \def\@@ylii{-513}\def\@@yli{-303}\fi\else
740 \ylposition{#2}{2}{5}{7}{0}%  %for 4-position
741 \fi

742 \if@ylsw\ifx\@@ylii\empty
743 \def\@@ylii{-513}\def\@@yli{-303}\fi\else
744 \ylfusedposition{#2}{-12}{-1}{1}{0}%  % for 4b and 8a
745 \fi
746 \if@ylsw\ifx\@@ylii\empty
747 \def\@@ylii{0}\def\@@yli{0}\fi\else
748 \ylfusedposition{#2}{-10}{0}{2}{0}% % for 4a and 10a
749 \fi
750 \if@ylsw\ifx\@@ylii\empty
751 \def\@@ylii{-342}\def\@@yli{0}\fi\fi
752 \if@ylsw
753 \yl@shiftii=\@ylii
754 \ifx\@@ylii\empty\else
755 \advance\yl@shiftii\@@ylii\fi
756 \yl@shifti=\@yli
757 \ifx\@@yli\empty\else
758 \advance\yl@shifti\@@yli\fi
759 \advance\yl@shiftii\yl@xdiff
760 \advance\yl@shifti\yl@ydiff
761 \begin{picture}(0,0)(-\yl@shiftii,-\yl@shifti)
762 \reset@yl@xydiff%1999/1/6 by S. Fujita
763 \else
764 \begin{picture}(1200,1200)(-\shiftii,-\shifti)
765 \iforigpt \put(-\shiftii,-\shifti){\circle*{50}}%
766  \put(-\noshift,-\noshift){\circle{50}}%
767 \typeout{command `hphenanthrenev' origin: %
768 (\the\noshift,\the\noshift) ---> (\the\shiftii,\the\shifti)}\fi%
769 \fi
770 \put(0,406){\line(-5,-3){171}}%       %bond 6-5
771 \put(0,406){\line(5,-3){171}}%        %bond 5-4a
772 \put(0,0){\line(-5,3){171}}%          %bond 8-7
773 \put(0,0){\line(5,3){171}}%            %bond 8-8a
774 \put(171,103){\line(0,1){200}}%       %bond 8a-4b
775 \put(-171,103){\line(0,1){200}}%      %bond 7-6
776 \put(342,406){\line(-5,-3){171}}%     %bond 4a-4b
777 \put(342,406){\line(5,-3){171}}%      %bond 4a-10a
778 \put(342,0){\line(-5,3){171}}%        %bond 9-8a
779 \put(342,0){\line(5,3){171}}%         %bond 9-10
780 \put(513,103){\line(0,1){200}}%       %bond 10-10a
781 %%%%
782 \put(342,406){\line(0,1){200}}%       %bond 4a-4
783 \put(513,709){\line(-5,-3){171}}%     %bond 3-4
784 \put(513,709){\line(5,-3){171}}%      %bond 3-2
785 \put(684,406){\line(0,1){200}}%       %bond 1-2
786 \put(513,303){\line(5,3){171}}%       %bond 10a-1
787 % inner double bonds
788 \@tfor\member:=#1\do{%
789 \if\member r%no endcyclic double bonds
790 \else\if\member a%
791 \put(651,432){\line(0,1){148}}%       %double bond 1-2

792 \else\if\member b%
793 \put(519,667){\line(5,-3){126}}%      %double bond 3-2
794 \else\if\member c%
795 \put(519,667){\line(-5,-3){126}}%     %double bond 3-4
796 \else\if\member d%
797 \put(375,432){\line(0,1){148}}%       %double bond 4a-4
798 \else\if\member e%
799 \put(348,364){\line(-5,-3){126}}%     %double bond 4a-4b
800 \else\if\member f%
801 \put(6,364){\line(5,-3){126}}%        %double bond 5-4b
802 \else\if\member g%
803 \put(-6,364){\line(-5,-3){126}}%      %double bond 5-6
804 \else\if\member h%
805 \put(-138,129){\line(0,1){148}}%      %double bond 6-7
806 \else\if\member i%
807 \put(-6,42){\line(-5,3){126}}%        %double bond 8-7
808 \else\if\member j%
809 \put(6,42){\line(5,3){126}}%          %double bond 8-8a
810 \else\if\member k%
811 \put(348,42){\line(-5,3){126}}%       %double bond 9-8a
812 \else\if\member l%
813 \put(348,42){\line(5,3){126}}%        %double bond 9-10
814 \else\if\member m%
815 \put(482,129){\line(0,1){148}}%       %double bond 10-10a
816 \else\if\member n%
817 \put(519,345){\line(5,3){126}}%       %double bond 10a-1
818 \else\if\member o%
819 \put(507,345){\line(-5,3){126}}%      %double bond 10a-4a
820 \else\if\member O%
821 \put(348,364){\line(5,-3){126}}%     %double bond 4a-10a
822 \else\if\member p%
823 \put(138,129){\line(0,1){148}}%       %double bond 8a-4b
824 \else\if\member P%
825 \put(204,129){\line(0,1){148}}%       %double bond 8a-4b
826 \else\if\member C%left aromatic circle
827 \put(0,203){\circle{240}}            %circle
828 \else\if\member B%right aromatic circle
829 \put(342,203){\circle{240}}%          %circle
830 \else\if\member A%right aromatic circle
831 \put(513,506){\circle{240}}%          %circle
832 \else
833 % setting fused rings
834 \expandafter\twoCH@R\member//%
835 \set@fusion@hphenanv
836 \fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi}%
837 %exocyclic bonds and substituents
838 \put(0,0){\setsixringv{#2}{-4}{0}{2}{0}}%   %for 5 position
839 \put(0,0){\setsixringv{#2}{12}{3}{7}{-1}}%   %for 6 to 8 positions
840 \put(342,0){\setsixringv{#2}{13}{2}{5}{-1}}%   %for 9 to 10 positions
841 \put(513,303){\setsixringv{#2}{4}{0}{4}{-1}}% %for 1 to 3 positions

842 \put(513,303){\setsixringv{#2}{2}{5}{7}{0}}%  %for 4-position
843 %bonds and substituents for fused positions
844 \put(0,0){\setfusedbond{#2}{-12}{-1}{1}{0}}%  % for 4b and 8a
845 \put(342,0){\setfusedbond{#2}{-10}{0}{2}{0}}% % for 4a and 10a
846 \end{picture}}%               %end of \hphenanthrenev macro%

10.2 Phenanthrene and phenanthroquinone derivatives

The macro \phenanthrenev is used for drawing phenanthrene derivatives of vertical type.

The macro \anthracenev has an argument SUBSLIST as well as an optional argument OPT. The latter designates a bond pattern in place of individual exocyclic bonds.

*********************************
* phenanthrene and  derivatives *
*  (vertical type)              *
*********************************
\phenanthrenev[OPT]{SUBSLIST}

The OPT argument contains one or two characters for designate aromatic bond patterns or for naphthoquinone bond patterns.

OPT (bond pattern)
none or r  :  right-handed double bonds
A          :  aromatic circle
p or pa    :  1,4-quinone (A)
pA         :  1,4-quinone (circle type)
o or oa    :  1,2-quinone (A)
oA         :  1,2-quinone (circle type)
ob         :  2,3-quinone (B)
oc         :  3,4-anthraquinone (C)
q or qa    :  9,10-quinone
qA         :  9,10-quinone (circle type)

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents

            nD         :  exocyclic double bond at n-atom
            n or nS    :  exocyclic single bond at n-atom
            nA         :  alpha single bond at n-atom
            nB         :  beta single bond at n-atom
            nSA        :  alpha single bond at n-atom (boldface)
            nSB        :  beta single bond at n-atom (dotted line)
            nSa        :  alpha (not specified) single bond at n-atom
            nSb        :  beta (not specified) single bond at n-atom
            nSd        :  alpha single bond at n-atom (dotted line)
                            with an alternative direction to nSA
            nSu        :  beta single bond at n-atom (boldface)
                            with an alternative direction to nSB
            nFA        :  alpha single bond at n-atom (dotted line)
                            for ring fusion
            nFB        :  beta single bond at n-atom (boldface)
                            for ring fusion
            nGA        :  alpha single bond at n-atom (dotted line)
                            for the other ring fusion
            nGB        :  beta single bond at n-atom (boldface)
                            for the other ring fusion

        e.g.
         \phenanthrenev{1==Cl;2==F}
         \phenanthrenev[A]{1==Cl;4==F;2==CH$_{3}$}
         \phenanthrenev[eb]{1D==O;4SA==MeO;4SB==OMe;%
                5W==Cl;6==Cl;7==Cl;8==Cl}

\phenanthrenev
847 \def\phenanthrenev{\@ifnextchar[{\@phenanthrenev}{\@phenanthrenev[r]}}
848 \def\@phenanthrenev[#1]#2{%
849 \iforigpt \typeout{command `phenanthrenev' is %
850  based on `hphenantherenv'.}\fi%
851 \expandafter\threech@r#1{}{}%
852 \if\@tmpa r% right-handed set of double bonds
853  \hphenanthrenev[acgilop]{#2}%
854 \else\if\@tmpa A%aromatic circle
855  \hphenanthrenev[ABC]{#2}%
856 \else\if\@tmpa p%1,4-quinone
857 \if\@tmpb a% (A)%
858  \hphenanthrenev[bgilop]{#2}%
859 \else\ifx\@tmpb\empty% (A)
860  \hphenanthrenev[bgilop]{#2}%
861 \else\if\@tmpb A% circle type
862  \hphenanthrenev[BCb]{#2}%

863 \else
864  \hphenanthrenev[bgilop]{#2}%
865 \fi\fi\fi%
866 \else\if\@tmpa o%o-quinone
867 \if\@tmpb a% (A)%
868  \hphenanthrenev[cgilop]{#2}%
869 \else\ifx\@tmpb\empty% (A)
870  \hphenanthrenev[cgilop]{#2}%
871 \else\if\@tmpb b% (B)
872  \hphenanthrenev[dgilnp]{#2}%
873 \else\if\@tmpb c% (C)
874  \hphenanthrenev[agilop]{#2}%
875 \else\if\@tmpb A% cicle type
876  \hphenanthrenev[cBC]{#2}%
877 \fi\fi\fi\fi\fi%
878 \else\if\@tmpa q%9,10-quinone
879 \if\@tmpb a% (A)%
880  \hphenanthrenev[acgiop]{#2}%
881 \else\ifx\@tmpb\empty% (A)
882  \hphenanthrenev[acgiop]{#2}%
883 \else\if\@tmpb A% circle type
884  \hphenanthrenev[AC]{#2}%
885 \else
886  \hphenanthrenev[acgiop]{#2}%
887 \fi\fi\fi%
888 \else
889  \hphenanthrenev[@#1]{#2}%
890 \fi\fi\fi\fi\fi%
891 }%               %end of \phenanthrenev macro%

11 Steroids

11.1 Steroids without a chain

The macro \steroid is used for drawing steroids.

The macro \steroid has an argument SUBSLIST as well as an optional argu-

 ment BONDLIST.
  ***********************
  * steroid derivatives *
  ***********************
    \steroid[BONDLIST]{SUBSLIST}

The BONDLIST argument contains one or more characters selected form a to t, each of which indicates the presence of an inner (endcyclic) double bond on the corresponding position. The character A, B, or C indicates an aromatic bond pattern (circle).

BONDLIST =
none:  steroid skeleton
a   :  1,2-double bond        b   :  2,3-double bond
c   :  3,4-double bond        d   :  4,5-double bond
e   :  6,5-double bond        f   :  6,7-double bond
g   :  7,8-double bond
h   :  9,8-double bond
h   :  9,8-double bond in the other bond
i   :  9,10-double bond       j   :  1,10-double bond
k   :  5,10-double bond
K   :  5,10-double bond in the other bond
l   :  9,11-double bond
m   :  12,11-double bond      n   :  12,13-double bond
o   :  14,13-double bond
O   :  14,13-double bond in the other bond
p   :  8,14-double bond
q   :  14,15-double bond      r   :  15,16-double bond
s   :  17,16-double bond      t   :  17,13-double bond
A   :  aromatic A ring        B   :  aromatic B ring
C   :  aromatic C ring

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 17 (execpt fused positions)
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom (boldface)
nB         :  beta single bond at n-atom (dotted line)
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion

            nGA        :  alpha single bond at n-atom (dotted line)
                            for the other ring fusion
            nGB        :  beta single bond at n-atom (boldface)
                            for the other ring fusion

For numbers larger than 9 (two digits), you should designate the SUBSLIST as

              {{11}A}==Cl;{{12}SA}==H;{{12}SB}==Cl;

        for n = 5,8,9,10,13 and 14 (fused positions)
            n or nS    :  exocyclic single bond at n-atom
            nA         :  alpha single bond at n-atom (boldface)
            nB         :  beta single bond at n-atom (dotted line)
            nU         :  unspecified single bond at n-atom

        e.g.
         \steroid{1==Cl;2==F;{{10}}==OH}
         \steroid[c]{1==Cl;4==F;2==CH$_{3}$;
             {{11}SA}==OH;{{11}SB}==OH}
         \steroid[eb]{1D==O;4SA==MeO;4SB==OMe;
                    5==Cl;6==Cl;7==Cl;8==Cl}

\steroid
\yl@steroidposition
\@steroidskeleton

892 \def\steroid{\@ifnextchar[{\@steroid[@}{\@steroid[Z]}}
893 \def\@steroid[#1]#2{%
894 % bonds for outer skeleton
895 \@reset@ylsw\reset@@yl
896 \yl@steroidposition{#2}%
897 \if@ylsw
898 \yl@shiftii=\@ylii
899 \ifx\@@ylii\empty\else
900 \advance\yl@shiftii\@@ylii\fi
901 \yl@shifti=\@yli
902 \ifx\@@yli\empty\else
903 \advance\yl@shifti\@@yli\fi
904 \advance\yl@shiftii\yl@xdiff
905 \advance\yl@shifti\yl@ydiff
906 \begin{picture}(0,0)(-\yl@shiftii,-\yl@shifti)
907 \reset@yl@xydiff%1999/1/6 by S. Fujita
908 \else
909 \begin{picture}(1600,1300)(-\shiftii,-\shifti)
910 \iforigpt \put(-\shiftii,-\shifti){\circle*{50}}%
911  \put(-\noshift,-\noshift){\circle{50}}%

912 \typeout{command `steroid' origin: %
913  (\the\noshift,\the\noshift) ---> (\the\shiftii,\the\shifti)}\fi%
914 \fi
915 \put(0,0){\@steroidskeleton{#1}{#2}}%
916 \end{picture}}%               %end of \steroid macro%
    The \yl@steroidposition is derived from \ylposition and \ylatombondposition

to use in the \steroid command as well as in the \steroidchain command.

917 \def\yl@steroidposition#1{%
918 \ylposition{#1}{0}{0}{2}{0}%  %for 1
919 \if@ylsw \ifx\@@ylii\empty
920 \def\@@ylii{0}\def\@@yli{0}\fi
921 \else
922 \ylposition{#1}{8}{3}{7}{-1}%  %for 2 to 4 positions
923 \fi
924 \if@ylsw\ifx\@@ylii\empty
925 \def\@@ylii{0}\def\@@yli{0}\fi\else
926 \ylposition{#1}{10}{2}{5}{-1}% %for 6 to 7 positions
927 \fi
928 \if@ylsw\ifx\@@ylii\empty
929 \def\@@ylii{-342}\def\@@yli{0}\fi\else
930 \ylposition{#1}{-11}{0}{2}{0}%for 12 position
931 \fi
932 \if@ylsw\ifx\@@ylii\empty
933 \def\@@ylii{-513}\def\@@yli{-303}\fi\else
934 \ylposition{#1}{-5}{5}{7}{0}% %for 11 position
935 \fi
936 \if@ylsw\ifx\@@ylii\empty
937 \def\@@ylii{-513}\def\@@yli{-303}\fi\else
938 \ylposition{#1}{18}{0}{4}{-1}%for 15 to 17 positions
939 \fi
940 \if@ylsw\ifx\@@ylii\empty
941 \def\@@ylii{-855}\def\@@yli{-303}\fi\else
942 \ylatombondposition{#1}{-4}{7}%     % for 5
943 \fi
944 \if@ylsw\ifx\@@ylii\empty
945 \def\@@ylii{-171}\def\@@yli{-103}\fi\else
946 \ylatombondposition{#1}{-9}{1}%     % for 10
947 \fi
948 \if@ylsw\ifx\@@ylii\empty
949 \def\@@ylii{-171}\def\@@yli{-303}\fi\else
950 \ylatombondposition{#1}{-7}{1}%     % for 8
951 \fi
952 \if@ylsw\ifx\@@ylii\empty
953 \def\@@ylii{-513}\def\@@yli{-303}\fi\else
954 \ylatombondposition{#1}{-8}{7}%     % for 9
955 \fi
956 \if@ylsw\ifx\@@ylii\empty
957 \def\@@ylii{-342}\def\@@yli{-406}\fi\else
958 \ylatombondposition{#1}{-13}{7}%    % for 14

959 \fi
960 \if@ylsw\ifx\@@ylii\empty
961 \def\@@ylii{-684}\def\@@yli{-406}\fi\else
962 \ylatombondposition{#1}{-12}{1}%    % for 13
963 \fi
964 \if@ylsw\ifx\@@ylii\empty
965 \def\@@ylii{-684}\def\@@yli{-606}\fi\fi%
966 }% end of \yl@steroidposition

The \@steroidskeleton is separated from the definition of \steroid so as to be used also in the definition of \steroidchain.

967 \def\@steroidskeleton#1#2{%
968 \put(0,406){\line(-5,-3){171}}%       %bond 1-2
969 \put(0,406){\line(5,-3){171}}%        %bond 1-10
970 \put(0,0){\line(-5,3){171}}%          %bond 4-3
971 \put(0,0){\line(5,3){171}}%           %bond 4-5
972 \put(171,103){\line(0,1){200}}%       %bond 5-10
973 \put(-171,103){\line(0,1){200}}%      %bond 3-2
974 \put(342,406){\line(-5,-3){171}}%     %bond 9-10
975 \put(342,406){\line(5,-3){171}}%      %bond 9-8
976 \put(342,0){\line(-5,3){171}}%        %bond 6-5
977 \put(342,0){\line(5,3){171}}%         %bond 6-7
978 \put(513,103){\line(0,1){200}}%       %bond 7-8
979 %%%%
980 \put(342,406){\line(0,1){200}}%       %bond 9-11
981 \put(513,709){\line(-5,-3){171}}%     %bond 12-11
982 \put(513,709){\line(5,-3){171}}%      %bond 12-13
983 \put(684,406){\line(0,1){200}}%       %bond 14-13
984 \put(513,303){\line(5,3){171}}%       %bond 8-14
985 %%%%
986 \put(855,709){\line(-5,-3){171}}%     %bond 17-13
987 \put(855,709){\line(5,-3){171}}%      %bond 17-16
988 \put(1026,406){\line(0,1){200}}%      %bond 15-16
989 \put(684,406){\line(1,0){342}}%       %bond 14-15
990 % inner double bonds
991 \@tfor\member:=#1\do{%
992 \if\member Z%no endcyclic double bonds
993 \else\if\member a%
994 \put(-6,364){\line(-5,-3){126}}%      %double bond 1-2 5-6
995 \else\if\member b%
996 \put(-138,129){\line(0,1){148}}%      %double bond 3-2 6-7
997 \else\if\member c%
998 \put(-6,42){\line(-5,3){126}}%        %double bond 4-3 8-7
999 \else\if\member d%
1000 \put(6,42){\line(5,3){126}}%          %double bond 4-5 8-8a
1001 \else\if\member e%
1002 \put(348,42){\line(-5,3){126}}%       %double bond 6-5 9-8a
1003 \else\if\member f%
1004 \put(348,42){\line(5,3){126}}%        %double bond 6-7 9-10
1005 \else\if\member g%

1006 \put(482,129){\line(0,1){148}}%       %double bond 7-8 10-10a
1007 \else\if\member h%
1008 \put(507,345){\line(-5,3){126}}%      %double bond 8-9  10a-4a
1009 \else\if\member H%
1010 \put(348,364){\line(5,-3){126}}%     %double bond 9-8  4a-10a
1011 \else\if\member i%
1012 \put(348,364){\line(-5,-3){126}}%     %double bond 9-10  4a-4b
1013 \else\if\member j%
1014 \put(6,364){\line(5,-3){126}}%        %double bond 1-10 5-4b
1015 \else\if\member k%
1016 \put(138,129){\line(0,1){148}}%       %double bond 5-10 8a-4b
1017 \else\if\member K%
1018 \put(204,129){\line(0,1){148}}%       %double bond 5-10 8a-4b
1019 \else\if\member l%
1020 \put(375,432){\line(0,1){148}}%       %double bond 9-11    4-5
1021 \else\if\member m%
1022 \put(519,667){\line(-5,-3){126}}%     %double bond 12-11   3-4
1023 \else\if\member n%
1024 \put(519,667){\line(5,-3){126}}%      %double bond 12-13   3-2
1025 \else\if\member o%
1026 \put(651,432){\line(0,1){148}}%       %double bond 13-14  1-2
1027 \else\if\member O%
1028 \put(717,432){\line(0,1){148}}%       %double bond 13-14  1-2
1029 \else\if\member p%
1030 \put(519,345){\line(5,3){126}}%       %double bond 8-14 10a-1
1031 \else\if\member q%
1032 \put(720,439){\line(1,0){260}}%       %double bond 14-15
1033 \else\if\member r%
1034 \put(993,432){\line(0,1){148}}%       %double bond 15-16  1-2
1035 \else\if\member s%
1036 \put(861,667){\line(5,-3){126}}%      %double bond 17-16   3-2
1037 \else\if\member t%
1038 \put(861,667){\line(-5,-3){126}}%     %double bond 17-13   3-4
1039 \else\if\member A%left aromatic circle
1040 \put(0,203){\circle{240}}%            %circle
1041 \else\if\member B%right aromatic circle
1042 \put(342,203){\circle{240}}%          %circle
1043 \else\if\member C%right aromatic circle
1044 \put(513,506){\circle{240}}%          %circle
1045 \else
1046 % setting fused rings
1047 \expandafter\twoCH@R\member//%
1048 \set@fusion@steroid
1049 \fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi%
1050 \fi\fi\fi\fi\fi\fi\fi\fi\fi}%
1051 %exocyclic bonds and substituents
1052 \iniflag
1053 \put(0,0){\setsixringv{#2}{0}{0}{2}{0}}%  %for 1
1054 \put(0,0){\setsixringv{#2}{8}{3}{7}{-1}}%  %for 2 to 4 positions
1055 \put(342,0){\setsixringv{#2}{10}{2}{5}{-1}}% %for 6 to 7 positions

1056 \put(513,303){\setsixringv{#2}{-11}{0}{2}{0}}%for 12 position
1057 \put(513,303){\setsixringv{#2}{-5}{5}{7}{0}}% %for 11 position
1058 \put(855,303){\setsixringv{#2}{18}{0}{4}{-1}}%for 15 to 17 positions
1059 %bonds and substituents for fused positions
1060 \put(171,103){\setatombond{#2}{-4}{7}}%     % for 5
1061 \put(171,303){\setatombond{#2}{-9}{1}}%     % for 10
1062 \put(513,303){\setatombond{#2}{-7}{1}}%     % for 8
1063 \put(342,406){\setatombond{#2}{-8}{7}}%     % for 9
1064 \put(684,406){\setatombond{#2}{-13}{7}}%    % for 14
1065 \put(684,606){\setatombond{#2}{-12}{1}}%    % for 13
1066 }% end of \@steroidskeleton

The macro \set@fusion@steroid is an inner command contained in the definition of \steroidskeleton, where the macro is used for setting an additive or a fused ring at each bond (designated by \@@tmpa). The fused ring is typeset by

  the inner command \set@fusionadd@sixv.
\set@fusion@steroid
1067 \def\set@fusion@steroid{%
1068 \let\@@tmpa=\@tmpa
1069 \let\@@tmpb=\@tmpb
1070 \if\@tmpa a%
1071  \edef\@@tmpa{f}%
1072  \put(0,0){\set@fusionadd@sixv}%
1073 \else\if\@tmpa A%
1074  \edef\@@tmpa{F}%
1075  \put(0,0){\set@fusionadd@sixv}%
1076 \else\if\@tmpa b%
1077  \edef\@@tmpa{e}%
1078  \put(0,0){\set@fusionadd@sixv}%
1079 \else\if\@tmpa B%
1080  \edef\@@tmpa{E}%
1081  \put(0,0){\set@fusionadd@sixv}%
1082 \else\if\@tmpa c%
1083  \edef\@@tmpa{d}%
1084  \put(0,0){\set@fusionadd@sixv}%
1085 \else\if\@tmpa C%
1086  \edef\@@tmpa{D}%
1087  \put(0,0){\set@fusionadd@sixv}%
1088 \else\if\@tmpa d%
1089  \edef\@@tmpa{c}%
1090  \put(0,0){\set@fusionadd@sixv}%
1091  \if@smallringsw\else
1092  \XyMTeXWarning{Unfavorable Fusion for bond `d'}%

1093 \fi

1094 \else\if\@tmpa D%
1095  \edef\@@tmpa{C}%
1096  \put(0,0){\set@fusionadd@sixv}%
1097  \if@smallringsw\else
1098  \XyMTeXWarning{Unfavorable Fusion for bond `D'}%

1099 \fi

1100 \else\if\@tmpa e%
1101  \edef\@@tmpa{d}%
1102  \put(342,0){\set@fusionadd@sixv}%
1103  \if@smallringsw\else
1104  \XyMTeXWarning{Unfavorable Fusion for bond `e'}%

1105 \fi

1106 \else\if\@tmpa E%
1107  \edef\@@tmpa{D}%
1108  \put(342,0){\set@fusionadd@sixv}%
1109  \if@smallringsw\else
1110  \XyMTeXWarning{Unfavorable Fusion for bond `E'}%

1111 \fi

1112 \else\if\@tmpa f%
1113  \edef\@@tmpa{c}%
1114  \put(342,0){\set@fusionadd@sixv}%
1115  \if@smallringsw\else
1116  \XyMTeXWarning{Unfavorable Fusion for bond `f'}%

1117 \fi

1118 \else\if\@tmpa F%
1119  \edef\@@tmpa{C}%
1120  \put(342,0){\set@fusionadd@sixv}%
1121  \if@smallringsw\else
1122  \XyMTeXWarning{Unfavorable Fusion for bond `f'}%

1123 \fi

1124 \else\if\@tmpa g%
1125  \edef\@@tmpa{b}%
1126  \put(342,0){\set@fusionadd@sixv}%
1127  \if@smallringsw\else
1128  \XyMTeXWarning{Unfavorable Fusion for bond `g'}%

1129 \fi

1130 \else\if\@tmpa G%
1131  \edef\@@tmpa{B}%
1132  \put(342,0){\set@fusionadd@sixv}%
1133  \if@smallringsw\else
1134  \XyMTeXWarning{Unfavorable Fusion for bond `g'}%

1135 \fi

1136 \else\if\@tmpa h%
1137  \edef\@@tmpa{a}%
1138  \put(342,0){\set@fusionadd@sixv}%
1139  \if@smallringsw\else
1140  \XyMTeXWarning{Forbidden Fusion for bond `h'}%

1141 \fi

1142 \else\if\@tmpa H%
1143  \edef\@@tmpa{A}%
1144  \put(342,0){\set@fusionadd@sixv}%
1145  \if@smallringsw\else
1146  \XyMTeXWarning{Forbidden Fusion for bond `H'}%

1147 \fi

1148 \else\if\@tmpa i%

1149  \edef\@@tmpa{f}%
1150  \put(342,0){\set@fusionadd@sixv}%
1151  \if@smallringsw\else
1152  \XyMTeXWarning{Unfavorable Fusion for bond `i'}%

1153 \fi

1154 \else\if\@tmpa I%
1155  \edef\@@tmpa{F}%
1156  \put(342,0){\set@fusionadd@sixv}%
1157  \if@smallringsw\else
1158  \XyMTeXWarning{Unfavorable Fusion for bond `I'}%

1159 \fi

1160 \else\if\@tmpa j%
1161  \edef\@@tmpa{a}%
1162  \put(0,0){\set@fusionadd@sixv}%
1163  \if@smallringsw\else
1164  \XyMTeXWarning{Unfavorable Fusion for bond `j'}%

1165 \fi

1166 \else\if\@tmpa J%
1167  \edef\@@tmpa{A}%
1168  \put(0,0){\set@fusionadd@sixv}%
1169  \if@smallringsw\else
1170  \XyMTeXWarning{Unfavorable Fusion for bond `J'}%

1171 \fi

1172 \else\if\@tmpa k%
1173  \edef\@@tmpa{b}%
1174  \put(0,0){\set@fusionadd@sixv}%
1175  \if@smallringsw\else
1176  \XyMTeXWarning{Forbidden Fusion for bond `k'}%

1177 \fi

1178 \else\if\@tmpa K%
1179  \edef\@@tmpa{B}%
1180  \put(0,0){\set@fusionadd@sixv}%
1181  \if@smallringsw\else
1182  \XyMTeXWarning{Forbidden Fusion for bond `K'}%

1183 \fi

1184 \else\if\@tmpa l%
1185  \edef\@@tmpa{e}%
1186  \put(513,303){\set@fusionadd@sixv}%
1187  \if@smallringsw\else
1188  \XyMTeXWarning{Unfavorable Fusion for bond `l'}%

1189 \fi

1190 \else\if\@tmpa L%
1191  \edef\@@tmpa{E}%
1192  \put(513,303){\set@fusionadd@sixv}%
1193  \if@smallringsw\else
1194  \XyMTeXWarning{Unfavorable Fusion for bond `L'}%

1195 \fi

1196 \else\if\@tmpa m%
1197  \edef\@@tmpa{f}%
1198  \put(513,303){\set@fusionadd@sixv}%

1199 \else\if\@tmpa M%
1200  \edef\@@tmpa{F}%
1201  \put(513,303){\set@fusionadd@sixv}%
1202 \else\if\@tmpa n%
1203  \edef\@@tmpa{a}%
1204  \put(513,303){\set@fusionadd@sixv}%
1205  \if@smallringsw\else
1206  \XyMTeXWarning{Unfavorable Fusion for bond `n'}%

1207 \fi

1208 \else\if\@tmpa N%
1209  \edef\@@tmpa{A}%
1210  \put(513,303){\set@fusionadd@sixv}%
1211  \if@smallringsw\else
1212  \XyMTeXWarning{Unfavorable Fusion for bond `N'}%

1213 \fi

1214 \else\if\@tmpa o%
1215  \edef\@@tmpa{b}%
1216  \put(513,303){\set@fusionadd@sixv}%
1217  \if@smallringsw\else
1218  \XyMTeXWarning{Forbidden Fusion for bond `o'}%

1219 \fi

1220 \else\if\@tmpa O%
1221  \edef\@@tmpa{B}%
1222  \put(513,303){\set@fusionadd@sixv}%
1223  \if@smallringsw\else
1224  \XyMTeXWarning{Forbidden Fusion for bond `O'}%

1225 \fi

1226 \else\if\@tmpa p%
1227  \edef\@@tmpa{c}%
1228  \put(513,303){\set@fusionadd@sixv}%
1229  \if@smallringsw\else
1230  \XyMTeXWarning{Unfavorable Fusion for bond `N'}%

1231 \fi

1232 \else\if\@tmpa P%
1233  \edef\@@tmpa{C}%
1234  \put(513,303){\set@fusionadd@sixv}%
1235  \if@smallringsw\else
1236  \XyMTeXWarning{Unfavorable Fusion for bond `N'}%

1237 \fi

1238 \else\if\@tmpa q%
1239  \XyMTeXWarning{Mismatched dimension for bond `q'}%
1240 \else\if\@tmpa Q%
1241  \XyMTeXWarning{Mismatched dimension for bond `Q'}%
1242 \else\if\@tmpa r%
1243  \edef\@@tmpa{b}%
1244  \put(855,303){\set@fusionadd@sixv}%
1245 \else\if\@tmpa R%
1246  \edef\@@tmpa{B}%
1247  \put(855,303){\set@fusionadd@sixv}%
1248 \else\if\@tmpa s%

1249  \edef\@@tmpa{a}%
1250  \put(855,303){\set@fusionadd@sixv}%
1251 \else\if\@tmpa S%
1252  \edef\@@tmpa{A}%
1253  \put(855,303){\set@fusionadd@sixv}%
1254 \else\if\@tmpa t%
1255  \edef\@@tmpa{f}%
1256  \put(855,303){\set@fusionadd@sixv}%
1257  \if@smallringsw\else
1258  \XyMTeXWarning{Unfavorable Fusion for bond `t'}%

1259 \fi

1260 \else\if\@tmpa T%
1261  \edef\@@tmpa{F}%
1262  \put(855,303){\set@fusionadd@sixv}%
1263  \if@smallringsw\else
1264  \XyMTeXWarning{Unfavorable Fusion for bond `T'}%

1265 \fi

1266 \fi\fi\fi\fi\fi\fi\fi\fi%
1267 \fi\fi\fi\fi\fi\fi\fi\fi\fi\fi%
1268 \fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi\fi%
1269 \fi\fi\fi\fi\fi\fi\fi\fi\fi\fi%
1270 }% end of the macro \set@fusion@steroid

11.2 Steroids with a chain

The macro \steroidchain is used for drawing steroids.

The macro \steroidchain has an argument SUBSLIST as well as an optional argument BONDLIST.

   **************************************
   * steroid derivatives having a chain *
   **************************************
     \steroidchain[BONDLIST]{SUBSLIST}

The BONDLIST argument contains one or more characters selected form a to t, each of which indicates the presence of an inner (endcyclic) double bond on the corresponding position. The character A, B, or C indicates an aromatic bond pattern (circle). A two-character string ‘Z?’ indicates a double bond in the sidechain.

       BONDLIST =
         (double bonds for the steroid skeleton)
             none:  no action on the steroid skeleton
             a   :  1,2-double bond        b   :  2,3-double bond
             c   :  3,4-double bond        d   :  4,5-double bond
             e   :  6,5-double bond        f   :  6,7-double bond

g   :  7,8-double bond        h   :  9,8-double bond
i   :  9,10-double bond       j   :  1,10-double bond
k   :  5,10-double bond       l   :  9,11-double bond
m   :  12,11-double bond      n   :  12,13-double bond
o   :  14,13-double bond      p   :  8,14-double bond
q   :  14,15-double bond      r   :  15,16-double bond
s   :  17,16-double bond      t   :  17,13-double bond
A   :  aromatic A ring        B   :  aromatic B ring
C   :  aromatic C ring

(double bonds for the side chain)
Z   :  no action
Za   :  17,20-double bond     Zb   :  20,22-double bond
Zc   :  22,23-double bond     Zd   :  23,24-double bond
Ze   :  24,25-double bond     Zf   :  25,26-double bond
Zg   :  25,27-double bond

For a two-character indicator, you should write the BONDLIST as, e.g.

   [b{Za}{Zc}...]

The SUBSLIST argument contains one or more substitution descriptors which are separated from each other by a semicolon. Each substitution descriptor has a locant number with a bond modifier and a substituent, where these are separated with a double equality symbol.

SUBSLIST: list of substituents
for n = 1 to 24 (execpt fused positions)
nD         :  exocyclic double bond at n-atom
n or nS    :  exocyclic single bond at n-atom
nA         :  alpha single bond at n-atom (boldface)
nB         :  beta single bond at n-atom (dotted line)
nSA        :  alpha single bond at n-atom (boldface)
nSB        :  beta single bond at n-atom (dotted line)
nSa        :  alpha (not specified) single bond at n-atom
nSb        :  beta (not specified) single bond at n-atom
nSd        :  alpha single bond at n-atom (dotted line)
                with an alternative direction to nSA
nSu        :  beta single bond at n-atom (boldface)
                with an alternative direction to nSB
nFA        :  alpha single bond at n-atom (dotted line)
                for ring fusion
nFB        :  beta single bond at n-atom (boldface)
                for ring fusion

             nGA        :  alpha single bond at n-atom (dotted line)
                             for the other ring fusion
             nGB        :  beta single bond at n-atom (boldface)
                             for the other ring fusion

For numbers larger than 9 (two digits), you should designate the SUBSLIST as

               {{11}A}==Cl;{{12}SA}==H;{{12}SB}==Cl; ...

         for n = 5,8,9,10,13,14,and 25 (fused positions etc.)
             n or nS    :  exocyclic single bond at n-atom
             nA         :  alpha single bond at n-atom (boldface)
             nB         :  beta single bond at n-atom (dotted line)
             nU         :  unspecified single bond at n-atom

         e.g.
          \steroidchain{1==Cl;2==F;{{10}}==OH}
          \steroidchain[c]{1==Cl;4==F;2==CH$_{3}$;%
              {{11}SA}==OH;{{11}SB}==OH}
          \steroidchain[b]{1D==O;4SA==MeO;4SB==OMe;%
                     ;6==Cl;7==Cl;11==Cl}

\steroidchain
1271 \def\steroidchain{\@ifnextchar[{\@steroidchain[@}{\@steroidchain[Z]}}
1272 \def\@steroidchain[#1]#2{%
1273 %shift due to the steroid skeleton
1274 \@reset@ylsw\reset@@yl
1275 \yl@steroidposition{#2}%
1276 %shift due to the side chain
1277 \ylposition{#2}{-21}{0}{3}{0}%for 22 and 23 positions
1278 \if@ylsw \ifx\@@ylii\empty
1279 \def\@@ylii{-1026}\def\@@yli{-606}\fi
1280 \else
1281 \ylposition{#2}{-14}{5}{7}{0}%for 20 position

1282 \fi

1283 \if@ylsw\ifx\@@ylii\empty
1284 \def\@@ylii{-1026}\def\@@yli{-606}\fi\else
1285 \ylposition{#2}{-20}{3}{5}{0}%for 24 position

1286 \fi

1287 \if@ylsw\ifx\@@ylii\empty
1288 \def\@@ylii{-1197}\def\@@yli{-709}\fi\else
1289 \ylatombondposition{#2}{-24}{4}%      % for 5

1290 \fi

1291 \if@ylsw\ifx\@@ylii\empty
1292 \def\@@ylii{-1397}\def\@@yli{-709}\fi\fi
1293 \if@ylsw
1294 \yl@shiftii=\@ylii
1295 \ifx\@@ylii\empty\else
1296 \advance\yl@shiftii\@@ylii\fi
1297 \yl@shifti=\@yli
1298 \ifx\@@yli\empty\else
1299 \advance\yl@shifti\@@yli\fi
1300 \advance\yl@shiftii\yl@xdiff
1301 \advance\yl@shifti\yl@ydiff
1302 \begin{picture}(0,0)(-\yl@shiftii,-\yl@shifti)
1303 \reset@yl@xydiff%1999/1/6 by S. Fujita
1304 \else
1305 \begin{picture}(2000,1500)(-\shiftii,-\shifti)
1306 \iforigpt \typeout{command `steroidchain' is based on `steroid'.}\fi%

1307 \fi

1308 % steroid skeleton
1309 \put(0,0){\@steroidskeleton{#1}{#2}}%
1310 % bonds for the side chain (carbons 20 to 27 (except 21))
1311 \put(855,709){\line(0,1){200}}%       %bond 17-20
1312 \put(1026,1012){\line(-5,-3){171}}%   %bond 22-20
1313 \put(1026,1012){\line(5,-3){171}}%    %bond 22-23
1314 \put(1197,709){\line(0,1){200}}%      %bond 24-23
1315 \put(1197,709){\line(1,0){200}}%      %bond 24-25
1316 \put(1397,709){\line(3,5){103}}%      %bond 25-26
1317 \put(1397,709){\line(3,-5){103}}%     %bond 25-27
1318 % double bonds along the side chain
1319 \@tfor\member:=#1\do{%
1320 \expandafter\twoCH@R\member//\relax%
1321 \if\@tmpa Z\relax%
1322 \ifx\@tmpb\empty%no action
1323 \else\if\@tmpb a%
1324 \put(888,735){\line(0,1){148}}%       %double bond 17-20
1325 \else\if\@tmpb b%
1326 \put(1020,970){\line(-5,-3){126}}%    %double bond 22-20
1327 \else\if\@tmpb c%
1328 \put(1032,970){\line(5,-3){126}}%     %double bond 22-23
1329 \else\if\@tmpb d%
1330 \put(1164,735){\line(0,1){148}}%      %double bond 24-23
1331 \else\if\@tmpb e%
1332 \put(1233,742){\line(1,0){140}}%      %double bond 24-25
1333 \else\if\@tmpb f%
1334 \put(1433,715){\line(3,5){72}}%       %double bond 25-26
1335 \else\if\@tmpb g%
1336 \put(1433,697){\line(3,-5){72}}%      %double bond 25-27
1337 \fi\fi\fi\fi\fi\fi\fi\fi\fi}
1338 % substitution on the side chain
1339 \put(1026,606){\setsixringv{#2}{-21}{0}{3}{0}}%for 22 and 23 positions
1340 \put(1026,606){\setsixringv{#2}{-14}{5}{7}{0}}%for 20 position

1341 \put(1197,709){\setsixringv{#2}{-20}{3}{5}{0}}%for 24 position
1342 \put(1397,709){\setatombond{#2}{-24}{4}}%      % for 5
1343 \end{picture}}%               %end of \steroidchain macro%

1344 ⟨/carom⟩